### 27 Feb 2024 v4

# Tight contact structures without symplectic fillings are everywhere

Abstract.
We show that for all $n\geq 3$, any $(2n+1)$-dimensional manifold that admits a tight contact structure, also admits a tight but non-fillable contact structure, in the same almost contact class. For $n=2$, we obtain the same result, provided that the first Chern class vanishes. We further construct Liouville but not Weinstein fillable contact structures on any Weinstein fillable contact manifold of dimension at least $7$ with torsion first Chern class.
1. Introduction
In contact topology, there is an important dichotomy dividing contact structures into those that are overtwisted and those that are tight. According to [Eli89, BEM15], the former are in fact topological and not geometric objects, in the sense that they satisfy $h$-principles saying that their existence and classification can be reduced to classical obstruction theory and algebraic topology. On the other hand, tight contact structures are inherently geometric, hence in general much harder to construct, and let alone classify. There is however an important sufficient condition for tightness: every contact manifold which is symplectically fillable [^1] Here we consider strong fillings, i.e. symplectic manifolds such that the contact structure at the boundary is defined by the contraction of the symplectic form with a Liouville vector field, locally defined near the boundary. is necessarily tight [Gro85, Eli90, Nie06, BEM15]. A natural question is then to understand the difference between the classes of tight and fillable contact structures. Our first result addresses this question in broad generality, on manifolds of dimension at least five.
Theorem A.
If $(M^{2n+1},\xi)$ with $n\geq 3$ is tight, then $M$ admits a tight, non-fillable contact structure in the same almost contact class. If $n=2$, the same holds, with the extra assumption that the first Chern class of $\xi$ is zero.
The assumption on the dimension in Theorem A is necessary, since in dimension three the statement is false. For example the sphere $\mathbb{S}^{3}$ admits a unique tight contact structure [Eli92], which is also symplectically fillable (the standard one). Historically the first examples of tight contact structures that are not symplectically fillable were found in dimension $3$ by Entyre-Honda [EH02] and on a special class of product manifolds by Massot-Niederkrüger-Wendl in dimensions 5 and greater [MNW13].
One can view Theorem A as suggesting that contact topology in higher dimensions exhibits significantly more flexibility than in dimension three and all contact phenomena ought to occur independent of the underlying smooth topology. In particular, the proof of Theorem A involves a reduction to the case of a standard smooth sphere, i.e. we first prove the following special case.
Theorem B.
For every $n\geq 2$, the sphere $\mathbb{S}^{2n+1}$ admits a tight, non-fillable contact structure that is homotopically standard.
Here, homotopically standard means homotopic to the standard contact structure among almost contact structures (the algebraic topological object underlying contact structures, in terms of which the $h$-principles in [Eli89, BEM15] are formulated). We then deduce Theorem A by considering connected sums with these non-fillable examples on spheres to preserve the smooth topology but alter the properties of the contact structure. Of course, this requires some significant effort since both the obstruction to fillability as well as tightness do not obviously behave well under connected sum.
Hierarchy of fillings
The notion of fillability comes with several different nuances, namely one has notions of Weinstein fillability, Liouville fillability and strong fillability. Thus ordered, each notion is stronger than the following one, and all of them in fact imply tightness. One thus has the following natural hierarchy of contact manifolds:
[…]
There has been a considerable amount of work over almost 30 years in studying these inclusions, and they are known to be strict in any odd dimension [Eli96, EH02, Ghi05, Gay06, NW11, Bow12, MNW13, BCS14, Zho21a, BGM22, GNE22]. On the other hand, all the examples known so far of contact structures with exotic fillability properties come from special geometric constructions, all of which again require the presence of non-trivial topology in the underlying smooth manifold. Now, as a consequence of Theorem A and Theorem B, the final inclusion is strict for all contact manifolds, as soon as the set of tight contact structures is non-empty.
Remark.
In fact, there is also the further notion of weak fillability, which as the name suggests is the weakest possible notion of fillability. Since this in fact agrees with strong fillability on spheres due to [MNW13, Proposition 6], we do not consider weak fillability here. In particular, even though strictly speaking in (the proof of) Theorem B we obstruct strong fillability, this equivalence of strong and weak fillability for spheres implies that those contact spheres in fact are also not weakly fillable.
Our next result addresses the strictness of the first inclusion in the hierarchy above by constructing examples that are Liouville fillable but admit no Weinstein filling on spheres of sufficiently large dimension.
Theorem C.
For any $n\geq 3$ there exist homotopically standard Liouville fillable contact structures on $\mathbb{S}^{2n+1}$ that are not Weinstein fillable.
By taking connected sum with examples on spheres we can further deduce that there are Liouville fillable contact manifolds without Weinstein fillings in the case that the first Chern class is torsion (cf. Theorem E below). Thus, other than addressing the $5$-dimensional case, it remains to determine the existence of strongly but not Liouville fillable contact structure on $\mathbb{S}^{2n+1}$, for all $n\geq 2$.
Infinite classes of examples
After one constructs examples of contact structures with exotic fillability properties, it is natural to try to produce infinitely many such examples as well. In this direction, by pushing further the techniques used for the proofs of the previous statements, namely this time taking contact connected sums with some other judiciously chosen Weinstein fillable contact structures on the sphere, we obtain the following.
Theorem D.
Let $n\geq 5$, and suppose that $M^{2n+1}$ admits a Weinstein fillable contact structure with torsion first Chern class. Then, $M$ admits infinitely many non-isomorphic, tight and not strongly fillable contact structures in the same almost contact class.
In particular, spheres of dimension at least $11$ admit infinitely many tight non-fillable contact structures which are homotopically standard. One also has a similar result for Liouville but not Weinstein fillable contact structures.
Theorem E.
Suppose that $M^{2n+1}$ is of dimension $2n+1\geq 7$ and admits a Weinstein fillable contact structure with torsion first Chern class. Then $M$ also admits infinitely many non-isomorphic Liouville fillable contact structures that are not Weinstein fillable in the same formal class.
Again, as a corollary, we obtain that spheres of dimension at least $7$ admit infinitely many Liouville but not Weinstein fillable contact structures which are homotopically standard.
Overview of proofs
In order to prove Theorem B we apply flexible contact surgery to examples that are known not to admit strong symplectic fillings, in order to obtain a (smoothly standard) sphere. Namely, we consider the Bourgeois construction on a sphere $\mathbb{S}^{2n-1}$ (with respect to a suitable open book), which yields a contact structure on $\mathbb{S}^{2n-1}\times\mathbb{T}^{2}$, and then kill the topology coming from the $\mathbb{T}^{2}$-factor via surgery. In doing so one has two major obstacles to deal with: namely one must ensure that tightness is preserved, and also that the obstructions to fillability persist. To ensure the former we use a dynamical criteria on the periodic orbits of suitable contact forms, which is a strong version of asymptotic dynamical convexity (ADC) introduced in [Laz20a, Zho21b], that in turn ensures algebraic tightness.
To find fillability obstructions that are not destroyed via surgery, we exploit ideas that have been used to classify fillings of such ADC-manifolds under stronger assumptions (such as symplectic asphericity, vanishing of $c_{1}$). This yields homological restrictions on such putative fillings, which can then be used themselves to provide obstructions. This approach was used for instance in [BGM22] to obstruct fillability of Bourgeois contact manifolds, where one used moduli spaces of closed spheres that appear after applying a certain capping procedure. We instead use a partial cap to build a strong cobordism to a flexibly fillable manifold, which admits a natural filling whose symplectic cohomology is known to vanish. One can then apply arguments classifying exact fillings of such flexible contact manifold as in [Zho22a], to obtain homological obstructions that can be made to persist under connected sum, thus also yielding Theorem A. In fact, a cobordism trick using [CE20, Laz20b] allows us to reduce the connected sum result to the case where the contact manifold we begin with is Weinstein fillable. In spite of this, the full proof does require that symplectic cohomology can be defined for general strong fillings, which has now been established using for example Pardon’s VFC package [Par19].
The proof of Theorem C yielding obstructions to Weinstein fillings are technically simpler, as one has exact fillings by assumption, and one can follow the uniqueness proofs without significant changes. Here the key step is to ensure that the surgery procedure can be applied to reach a standard sphere in the first place.
Outline of Paper
Section 2 contains important preliminaries and background, including a description of the Bourgeois construction together with a result concerning dynamical convexity of its Reeb vector field, as well as statements about the existence of strong cobordisms from contact manifolds, obtained from subcritical (resp. flexible) surgeries on $\mathbb{S}^{1}-$equivariant ones, to subcritically (resp. flexibly) fillable ones.
In Section 3, we describe then the notion of algebraic tightness, together with relevant properties, and prove that the Bourgeois construction is a rich source of contact manifolds satisfying this property.
Section 4 contains results on obstructions to strong fillability via symplectic cohomology and Section 5 contains the proof of Theorem B and Theorems A and D. Lastly, Section 6 contains the proofs of Theorems C and E.
Acknowledgements
A. Moreno received support from the National Science Foundation under Grant No. DMS-1926686, and is currently supported by the Sonderforschungsbereich TRR 191 Symplectic Structures in Geometry, Algebra and Dynamics, funded by the DFG (Projektnummer 281071066 – TRR 191), and also by the DFG under Germany’s Excellence Strategy EXC 2181/1 - 390900948 (the Heidelberg STRUCTURES Excellence Cluster). Z. Zhou is supported by National Natural Science Foundation of China under Grant No. 12288201 and 12231010. F. Gironella gratefully acknowledges the support of the Sonderforschungsbereich Higher Invariants 1085. This project has also received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 772479), as well as from the Agence Nationale de la Recherche (ANR) under the Project COSY (ANR-21-CE40-0002-04). The authors are moreover grateful to Peter Albers and Yasha Eliashberg, for productive discussions and interest in the project, and to Frédéric Bourgeois, Roger Casals, Otto van Koert, Oleg Lazarev, Alex Oancea, for helpful feedback on an earlier version of this paper.
2. Bourgeois contact manifolds
In this section, we recall the Bourgeois construction [Bou02b], following the presentation that was given in [BGM22].
Consider a closed, oriented, connected smooth manifold $M^{2n-1}$ and an open book decomposition that we denote $(B,\theta)$, together with a defining map $\Phi\colon\thinspace M\rightarrow\mathbb{R}^{2}$ so that each $z\in\operatorname{int}(\mathbb{D}^{2})$ is a regular value. Here, $B\subset M$ is a closed codimension-2 submanifold, $\theta\colon M\backslash B\rightarrow\mathbb{S}^{1}$ is a fiber bundle, and $\Phi$ is such that $\Phi^{-1}(0)=B$ and $\theta=\Phi/\left|\Phi\right|$. Let us also denote by $\rho$ the norm $|\Phi|$.
Recall that a $1$-form $\alpha$ on $M$ is said to be adapted to $\Phi$ if it induces a contact structure on the regular fibers of $\Phi$ and if $d\alpha$ is symplectic on the fibers of $\theta=\Phi/\left|\Phi\right|$. In particular, if $\xi$ is a contact structure on $M$ supported by $(B,\theta)$, in the sense of [Gir02], then there is such a pair $(\alpha,\Phi)$ with $\alpha$ defining $\xi$ as follows from the definition.
Theorem 2.1 (Bourgeois [Bou02b]).
Let $(B,\theta)$ denote an open book decomposition of $M^{2n-1}$, represented by a map $\Phi=(\Phi_{1},\Phi_{2})\colon\thinspace M\rightarrow\mathbb{R}^{2}$ as above, and let $\alpha$ be a $1$-form adapted to $\Phi$. Then, $\alpha_{BO}\coloneqq\thinspace\alpha+\Phi_{1}dq_{1}-\Phi_{2}dq_{2}$ is a contact form on $M\times\mathbb{T}^{2}$, where $(q_{1},q_{2})$ are coordinates on $\mathbb{T}^{2}$.
The contact form $\alpha_{BO}$ on $M\times\mathbb{T}^{2}$ will be called the Bourgeois form associated to $(\alpha,\Phi)$ in what follows. It will be useful to also think in terms of abstract open books, i.e. in terms of the Liouville page $\Sigma=(\Sigma,d\lambda)$ and the compactly supported symplectic monodromy $\psi$. In this case we write $OB(\Sigma,\psi)$ for the contact manifold obtained via the Thurston-Winkelnkemper construction, and we denote $BO(\Sigma,\psi)=(M\times\mathbb{T}^{2},\ker\alpha_{BO})$. We invite the reader to consult [LMN19] for details on the correspondence between Bourgeois contact structures for geometric and abstract open books.
Remark 2.2.
The almost contact structure underlying the Bourgeois contact manifold is (up to homotopy) just the sum
$\left(\xi,d\alpha\right)\oplus(T\mathbb{T}^{2},\Omega),$
where $\Omega$ is an area form on $\mathbb{T}^{2}$. This can be seen via an explicit homotopy; see for instance [Gir20a, Lemma 4.1.(1)]. In particular if $c_{1}(\xi)$ is torsion, then the same is true for the Bourgeois contact structure, independently of the auxiliary choice of open book.
2.1. Reeb dynamics of Bourgeois contact structures
In this somewhat technical section, we consider the Reeb dynamics of the Bourgeois contact forms more closely. First of all applying a lemma of Giroux (see e.g. [DGZ14, Section 3]), on a neighborhood of the form $\mathcal{N}=B\times\mathbb{D}^{2}\times\mathbb{T}^{2}$ of $B$ where $\rho=|\Phi|$ coincides with the radial coordinate of $\mathbb{D}^{2}$, the Bourgeois contact form looks like
$\alpha_{BO}$ $=\alpha+\rho(r)(\cos(\theta)dq_{1}-\sin(\theta)dq_{2})$
$=h_{1}(r)\alpha_{B}+h_{2}(r)d\theta+r(\cos(\theta)dq_{1}-\sin(\theta)dq_{2}).$
Here, $\alpha_{B}=\alpha|_{B}$ is the contact form on the binding, $(r,\theta)$ are polar coordinates on $\mathbb{D}^{2}$, and the functions $h_{1},h_{2}$ satisfy the following conditions:
  1. i.
    $h_{1}(0)>0$ and $h_{1}(r)=h_{1}(0)+O(r^{2})$ for $r\rightarrow 0$;
  2. ii.
    $h_{2}(r)\sim r^{2}$ for $r\rightarrow 0$;
  3. iii.
    $\frac{h_{1}^{n-1}}{r}\,(h_{1}h_{2}^{\prime}-h_{2}h_{1}^{\prime})>0$ for $r\geq 0$ (contact condition);
  4. iv.
    $h_{1}^{\prime}(r)<0$ for $r>0$ (symplectic condition on the pages).
We also point out that there is a natural (orientation preserving) diffeomorphism
$\begin{split}\mathcal{N}=B\times\mathbb{D}^{2}\times\mathbb{T}^{2}&\to B\times D^{*}\mathbb{T}^{2}\\ (b,p_{1},p_{2},q_{1},q_{2})&\mapsto(b,p_{1},q_{1},-p_{2},q_{2}).\end{split}$ (1)
We then have the following global description of the Reeb vector field (c.f. [Bou02b, Gir20a]):
Lemma 2.3.
The Reeb vector field of the contact form $\alpha_{BO}$ is given by
$R_{BO}=\mu(r)R_{B}+\nu(r)[\sin(\theta)\partial_{q_{1}}-\cos(\theta)\partial_{q_{2}}],$ (2)
where $R_{B}$ is the Reeb vector field of the restriction $\alpha_{B}=\alpha|_{B}$, and the coefficients are
$\mu\,=\,\begin{cases}\frac{\rho^{\prime}}{\rho^{\prime}h_{1}-\rho h_{1}^{\prime}}&\text{in }\mathcal{N}\\ 0&\text{elsewhere }\end{cases}\,\,\text{ and }\,\,\nu\,=\begin{cases}\frac{-h_{1}^{\prime}}{\rho^{\prime}h_{1}-\rho h_{1}^{\prime}}&\text{in }\mathcal{N}\\ 1&\text{elsewhere }\end{cases}\text{ .}$
The computation in [Bou02b] shows that, in order for $\alpha_{BO}$ to be a contact form in the neighborhood $\mathcal{N}\simeq B\times D^{*}\mathbb{T}^{2}$ of $B\times\mathbb{T}^{2}$ where $\alpha=h_{1}\alpha_{B}+h_{2}d\theta$, it is in fact enough that $h_{1}(h_{1}-h_{1}^{\prime})>0$, a condition which only depends on $h_{1}$. In particular, one can homotope the pair $(h_{1},h_{2})$ in a compactly supported way among pairs of functions satisfying this condition, and this will result in a homotopy of contact forms (hence isotopy by Gray’s stability) on $M\times\mathbb{T}^{2}$, independently of the fact that the resulting $\alpha$ on $M$ might not be adapted to the open book or even a contact form. Moreover, the explicit formula in Lemma 2.3 still holds for the homotoped contact form, as the explicit computation does not use any specific property of the pair of functions $(h_{1},h_{2})$ listed above. In particular, up to homotopy one can achieve the following form, which will be useful below:
Lemma 2.4.
For $\delta>0$ sufficiently small, up to a deformation among contact structures on $M\times\mathbb{T}^{2}$ supported in the neighbourhood of radius $2\delta$ of $B\times\mathbb{T}^{2}$, one can assume that
  • $h_{1}(r)=1$ for $r\leq\delta$,
  • $h_{1}(h_{1}-h_{1}^{\prime})>0$ everywhere,
  • $h_{2}(r)=0$ for $r\leq 3\delta/2$.
In particular, under the diffeomorphism in Equation 1, $\alpha_{BO}$ coincides with $\alpha_{B}+\lambda_{std}$ for $r=\sqrt{p_{1}^{2}+p_{2}^{2}}\leq\delta$, where $\lambda_{std}=p_{1}dq_{1}+p_{2}dq_{2}$ is the standard Liouville form on $D^{*}\mathbb{T}^{2}$. Moreover, Equation 2 also holds for the Reeb vector field of the deformed contact form; in particular, it coincides with $R_{B}$ for $r\leq\delta$.
2.2. ADC Bourgeois contact manifolds
In this section, we will assume that the first Chern class $c_{1}(\xi)$ of all contact structures are torsion. In this case there is a well-defined (rational) Conley-Zehnder index $\mu_{CZ}(\gamma)$ for any non-degenerate contractible periodic Reeb orbit $\gamma$. If the contact manifold is of dimension $12n-$ 1, then the SFT degree of $\gamma$ is defined as $|\gamma|\coloneqq n-3+\mu_{CZ}(\gamma)$. Notice that this does not correspond to the degree in symplectic cohomology of $\gamma$ when viewed as a periodic Hamiltonian orbit, which is just given by $n$ minus the Conley-Zehnder index.
We now recall the following definition from [Zho21b], that generalizes the one from [Laz20a]:
Definition 2.5.
A contact structure $(M,\xi)$ is called $k$ asymptotically dynamically convex ($k$-ADC) if there is a sequence of contact forms $\alpha_{1}\geq\alpha_{2}\geq\cdots\geq\alpha_{i}\geq\dots$ so that all contractible periodic orbits of $\alpha_{i}-$action $\leq D_{i}$ are non-degenerate and have degree $>k$, with $D_{i}\to\infty$.
In particular, 00-ADC as in [Zho21b] is just ADC as in [Laz20a]. An important special case is when there is a contact form $\alpha$ that is index-positive, meaning, following [CO18, Section 9.5], that all contractible periodic orbits are non-degenerate and have positive degree for $\alpha_{i}\equiv\alpha$ the given contact form. A very useful aspect of the ADC condition is that it is preserved under flexible surgery [Laz20a], whereas index-positivity is not.
Example 2.6 (Brieskorn Manifolds).
Consider the $12n-$ 1-dimensional Brieskorn spheres $\Sigma_{n}(k,2,\cdots,2)$ given as the link of the $A_{k-1}$-singularity, i.e.
$\Sigma_{n}(k,2,\cdots,2)=\{z_{0}^{k}+z_{1}^{2}+\cdots+z_{n}^{2}=0\}\cap S^{2n+1}\subset\mathbb{C}^{n+1}.$
This is a contact manifold which is index-positive. The degrees of the generators are at least $42n-$ 4, provided that $n\geq 3$. This follows from [Ust99] in case $n=2m+1$ is odd, and from [vK08, Section 3] in general.
More generally, one can consider Brieskorn spheres
$\Sigma_{n}(p_{1},p_{2},\cdots,p_{n-2},2,2,2)$
for odd primes $p_{i}$, which according to [vK08, Section 4.1], satisfy the same properties in terms of their Conley-Zehnder indices. This also follows from a deep theorem of McLean [McL16, Theorem 1.1], which implies that the Brieskorn manifold $\Sigma_{n}(p_{1},p_{2},\cdots,p_{n+1})$ is ADC if the singularity is terminal.
Example 2.7 (3-dimensional case).
For the case of $3$-dimensional Brieskorn manifolds, the computations are more involved as the manifolds are not simply connected. One can view links of $A_{k-1}$-singularities as quotients of the standard contact $3$-sphere so that contractible periodic orbits on the links can be lifted to the $3$-sphere, which then have Conley-Zehnder indices at least $3$. However, when we construct a $7$-dimensional Bourgeois contact structure $\mathbb{S}^{5}\times\mathbb{T}^{2}$ with the $3$-dimensional $A_{k-1}$-singularity link as the binding, these non-contractible orbits in the binding become contractible in the Bourgeois contact manifold. Since $H_{1}(\Sigma(k,2,2);\mathbb{Z})$ is torsion, all Reeb orbits, including those non-contractible orbits, can be assigned with well-defined rational Conley-Zehnder indices [McL16, §4], which are at least $1$ by the computation in [MR18, §2.1]. In view of this, the 3-dimensional links of $A_{k-1}$-singularities are $-1$-ADC.
We next observe that the $k$-ADC condition is stable under the Bourgeois construction:
Lemma 2.8.
If the binding of an open book $OB(\Sigma,\psi)$ is simply connected and $k$-ADC, then the Bourgeois contact manifold $BO(\Sigma,\psi)$ is $(k+2)$-ADC.
Proof.
In view of Lemma 2.4, we can assume that, for some fixed $\delta>0$ sufficiently small, on a neighborhood $B\times\mathbb{D}^{2}_{2\delta}\times\mathbb{T}^{2}$ of the (stabilized) binding of the open book, the Bourgeois contact form is
$\alpha_{BO}=h_{1}(r)\alpha_{B}+h_{2}(r)d\theta+\lambda_{std}$
where $\lambda_{std}$ is the standard Liouville form on the portion $\mathbb{D}^{2}_{2\delta}\times\mathbb{T}^{2}\simeq D^{*}_{2\delta}\mathbb{T}^{2}$ of unit cotangent bundle made of vectors of norm less than $2\delta$, and where the functions $h_{1},h_{2}$ are such that the following conditions hold:
  • $h_{1}(r)\equiv 1$ for $r\leq\delta$ and $h_{1}^{\prime}(r)<0$ otherwise;
  • $h_{2}(r)\equiv 0$ for $r\leq 3\delta/2$ and $h_{2}^{\prime}(r)>0$ otherwise.
Consider then the Hamiltonian $H(q_{1},p_{1},q_{2},p_{2})=\eta(r)(f(q_{1},q_{2})+g(r)),$ where $(q_{i},p_{i})$ are coordinates on $D_{2\delta}^{*}\mathbb{T}^{2}$, $r=\sqrt{p_{1}^{2}+p_{2}^{2}}$, $\eta$ is a bump cut-off function, $f\colon\mathbb{T}^{2}\to\mathbb{R}$ is a Morse-function on $\mathbb{T}^{2}$, and $g\colon[0,1]\rightarrow\mathbb{R}$ is Morse with a unique maximum at $r=0$. Following Bourgeois [Bou02a] we consider
$\alpha_{\epsilon}=(h_{1}+\varepsilon H)\alpha_{B}+\lambda_{std}$
on the $2\delta$-neighbourhood $B\times\mathbb{D}^{2}_{2\delta}\times\mathbb{T}^{2}$. Then one computes that the Reeb vector field is of the form
$\frac{1}{h_{1}+\varepsilon H-\lambda_{std}(X_{h_{1}}+\varepsilon X_{H})}\left[R_{B}-X_{h_{1}}-\varepsilon X_{H}\right],$
where we use the notation $X_{h}$ for the Hamiltonian vector field of a function $h:D^{*}\mathbb{T}^{2}\rightarrow\mathbb{R}$ with respect to $\omega_{std}=d\lambda_{std}$. Note that the flow generated by the vector field $X_{h_{1}}$ is just a rescaling of the geodesic flow on $D^{*}\mathbb{T}^{2}$ and is stationary precisely where $h_{1}^{\prime}=0$.
Now for a fixed action threshold $D$, taking $\varepsilon$ small, we see that the only contractible periodic orbits up to action $D$ must lie in the region where $h_{1}$ is constant, i.e. $B\times\mathbb{D}^{2}_{\delta}\times\mathbb{T}^{2}$. These orbits are of the form $\gamma_{q}=\gamma_{B}\times\{(q,r=0)\}$ with $q$ a critical point of $f$ and $\gamma_{B}$ a contractible Reeb orbit in the binding. Moreover, one can compute, using the properties of the Conley-Zehnder index (c.f. [Sal99, Section 2.4]) that
$\mu_{CZ}(\gamma_{q})=\mu_{CZ}(\gamma_{B})-2+\textrm{ind}_{q}(H)\geq\mu_{CZ}(\gamma_{B}),$
as $\mathrm{ind}_{q}(H)=\mathrm{ind}_{q}(f)+2\geq 2$. Thus the degree on $M^{2n-1}\times\mathbb{T}^{2}$ is
$|\gamma_{q}|=(n+1)-3+\mu_{CZ}(\gamma_{q})\geq(n-1)-3+\mu_{CZ}(\gamma_{B})+2=|\gamma_{B}|_{B}+2,$
where $|\gamma_{B}|_{B}$ denotes the degree of $\gamma_{B}$ as an orbit in $B$. Consider a decreasing sequence $\alpha_{B}^{i}$ on the binding whose contractible orbits have degrees at least $k$ as in the definition of $k$-ADC, i.e. up to some action threshold $D_{i}\to\infty$. By multiplying by appropriate constants we can assume that this sequence is strictly decreasing. We then multiply the adapted contact form on $OB(\Sigma,\psi)$ by an appropriate function $f_{i}<1$ so that the form on the binding $\alpha_{B}$ is replaced by $\alpha_{B}^{i}$ to obtain a decreasing sequence of forms $\alpha^{i}$ for the Bourgeois contact structure whose contractible orbits also have degrees at least $k+2$ up to the action threshold $D_{i}$. Then perturbing as above to obtain non-degeneracy near the binding, we obtain contact forms $\alpha_{\varepsilon_{i}}^{i}$ so that the action of the above contractible orbit $\gamma_{q}$ is $(1+\varepsilon_{i}H(q,0))$ times its $\alpha_{B}^{i}$-action and $\varepsilon_{i}\ll 1$ small.
Strictly speaking these contact forms determine distinct contact structures, as their kernels may differ, but applying Gray stability this can be remedied by a isotopy, which can be made arbitrarily close to the identity by shrinking $\varepsilon_{i}$ as needed. We deduce that the Bourgeois contact structure is $(k+2)$-ADC if the contact structure on $B$ is $k$-ADC. ∎
Remark 2.9.
When $B$ is not simply connected, there might be Reeb orbits of $B$ which are non-contractible in $B$ but contractible in the ambient open book $OB(\Sigma,\psi)$. These orbits should then be included in the verification of the ADC condition for the open book, even though they are not relevant for that of $B$. In our applications, this will only happen in settings where $H_{1}(B;\mathbb{Z})$ is torsion. In this case, if the contact structure on $B$ has torsion first Chern class, every Reeb orbit can be assigned with a well-defined rational Conley-Zehnder index [McL16, §4]. One can then consider the notion of $k$-ADC manifolds but for all Reeb orbits. The computation in Lemma 2.8 is still valid for the rational Conley-Zehnder index.
We have the following examples that will be used in the proof of Theorem B.
Example 2.10.
Consider the Bourgeois contact manifold associated to the Milnor open book on $\mathbb{S}^{2n-1}$ coming from the $A_{k-1}$-singularity, whose binding is a Brieskorn sphere $B=\Sigma_{n-1}(k,2,\cdots,2)$ and whose page is the Milnor fiber $V_{n-1}(k,2,\cdots,2)=\{z_{0}^{k}+z_{1}^{2}+\dots+z_{n-1}^{2}=\epsilon\}\cap\mathbb{D}^{2n}$, for small $\epsilon>0$. When $n\geq 4$, $B$ is a $1$-ADC contact manifold according to Example 2.6 and simply connected. We thus obtain from Lemma 2.8 that the corresponding Bourgeois contact manifolds on $\mathbb{S}^{2n-1}\times\mathbb{T}^{2}$ are $1$-ADC. When $n=3$, the corresponding Bourgeois contact manifolds on $\mathbb{S}^{2n-1}\times\mathbb{T}^{2}$ are also $1$-ADC by Example 2.7 and Remark 2.9.
2.3. $\mathbb{S}^{1}-$invariant contact structures and convex decompositions
Bourgeois contact structures are $\mathbb{T}^{2}$-invariant, and therefore in particular $\mathbb{S}^{1}$-invariant. As discussed in [DG12, Section 6] any $\mathbb{S}^{1}-$invariant contact structure on a manifold $V^{2n}\times\mathbb{S}^{1}$ induces a decomposition of the first factor into so called ideal Liouville domains, as defined in [Gir20b]. In particular, we have a topological decomposition
$V\times\mathbb{S}^{1}=V_{+}\times\mathbb{S}^{1}\bigcup\overline{V}_{-}\times\mathbb{S}^{1}.$
Any (constant) section $V\times\{pt\}$ is a convex hypersurface and the decomposition $V=V_{+}\cup\overline{V}_{-}$ is a convex decomposition. In the case of the Bourgeois contact manifolds the ideal Liouville pieces are given as products of the page of the initial open book with an annulus, i.e. $V_{\pm}=\Sigma\times\mathbb{D}^{*}\mathbb{S}^{1}$. For details related to the notion of ideal Liouville domains and to the Bourgeois case we refer to [DG12, MNW13, Gir20b]. The pieces in this decomposition are round contactizations of ideal Liouville domains and are referred to as Giroux domains following [MNW13].
2.4. Capping Giroux domains
In [BGM22, Theorem B] certain homological obstructions to the symplectic fillability of Bourgeois contact manifolds were obtained. The argument of the proof uses a capping construction of Massot-Niederkrüger-Wendl [MNW13]. More precisely, whenever one finds an embedded Giroux domain of the form $G=X\times\mathbb{S}^{1}$, one can perform a blow down construction as in [MNW13], which topologically consists in removing the interior of $G$ and collapsing all circle fibers at the remaining boundary to points. By [MNW13, Section 5.1] the resulting smooth manifold $M_{bd}$ (where $bd$ stands for blown down) carries a contact structure that is well-defined (up to isotopy). The following special case of [MNW13, Theorem 6.1] then says that this is achievable through a cobordism:
Theorem 2.11.
Let $M$ be a contact manifold containing a Giroux domain $G=X\times\mathbb{S}^{1}$. Then there is a strong symplectic cobordism from $M$ to the contact manifold $M_{bd}$ obtained by blow down of $G$.
Topologically, this cobordism is obtained by attaching $X\times\mathbb{D}^{2}$ on top of $X\times\mathbb{S}^{1}$ (and smoothing corners), where the latter is seen as the positive boundary of $[0,1]\times X\times\mathbb{S}^{1}$. We emphasize that this cobordism is not Liouville, as there are only local Liouville vector fields near the boundary in general.
In the case where we begin with an $\mathbb{S}^{1}$-invariant contact structure and apply this blow down cobordism to Giroux domain $V_{-}\times\mathbb{S}^{1}$, we end up with a convex boundary $M_{bd}$ which is just the contact open book $OB(V_{+},\operatorname{Id})$ with page $V_{+}$ and trivial monodromy. Equivalently this is then the contact boundary of $V_{+}\times\mathbb{D}^{2}$. In other words, we have the following:
Corollary 2.12.
Let $M=V\times\mathbb{S}^{1}$ be a contact manifold be a contact manifold with an $\mathbb{S}^{1}$-invariant contact structure and convex decomposition
$V\times\mathbb{S}^{1}=V_{+}\times\mathbb{S}^{1}\bigcup\overline{V}_{-}\times\mathbb{S}^{1}.$
Then there is a strong symplectic cobordism from $M$ to the contact boundary of $V_{+}\times\mathbb{D}^{2}$. Furthermore, if $c_{1}(M)$ is torsion, then the same is true for the cobordism.
2.5. Allowing surgeries on a Giroux domain
We now wish to state a generalisation of Corollary 2.12, where we allow contact surgeries in one part of the convex decomposition.
Proposition 2.13.
Consider an $\mathbb{S}^{1}$-equivariant contact structure on $M=V^{2n}\times\mathbb{S}^{1}$, with induced convex splitting $V=V_{+}\cup\overline{V}_{-}$. Suppose now that $(M^{\prime},\xi^{\prime})$ is obtained from the former via a sequence of subcritical/flexible surgeries on isotropic/loose Legendrian spheres contained in the complement of $V_{-}\times\mathbb{S}^{1}$.
Then there is a strong symplectic cobordism $W_{cap}$ from $(M^{\prime},\xi^{\prime})$ to the contact manifold $(M_{+}^{\prime},\xi_{+}^{\prime})$ obtained from $OB(V_{+},\operatorname{Id})$ via the corresponding sequence of contact surgeries (i.e. those same contact surgeries performed on $V_{+}\times\mathbb{S}^{1}$ considered as a subset of $OB(V_{+},\operatorname{Id})$).
Moreover, if $V_{+}$ is Weinstein, there is a (smooth) copy of $V_{+}$ inside $M_{+}^{\prime}$, such that the following holds:
  • $V_{+}\subset M_{+}^{\prime}$ can be isotoped to lie in the negative end $M^{\prime}$ of the cobordism $W_{cap}$;
[…]
Proof.
The existence of the cobordism simply follows from the fact that the blow-down cobordism of Theorem 2.11 can be described as an attachment of a handle $V_{-}\times\mathbb{D}^{2}$ on top of $V_{-}\times\mathbb{S}^{1}\subset M$, and that the contact surgeries are made in the interior of $V_{+}\times\mathbb{S}^{1}\subset M$, which is disjoint from $V_{-}\times\mathbb{S}^{1}$; see Figure 1. In particular, in the symplectic cobordism given by blowing-down and performing the contact surgeries, one can find a symplectic sub-cobordism from $M^{\prime}$ to $M_{+}^{\prime}$ as claimed in the statement.
For the second part of the statement, first notice that, because of the dimensions involved, the attaching regions of the surgery handles that give $M_{+}^{\prime}$ starting from $OB(V_{+},\operatorname{Id})$ can be made smoothly disjoint from the $n$-skeleton $V_{+}^{(n)}$ of a page $V_{+}$ in the open book by general position. In particular, by the Weinstein assumption on $V_{+}$, one naturally has a copy of $N(V_{+}^{(n)})\simeq V_{+}$ inside $M_{+}^{\prime}$. What is more, as $M_{+}^{\prime}$ is topologically obtained from $M^{\prime}$ by attaching a handle $V_{-}\times\mathbb{D}^{2}$, i.e. Morse handles of indices between $1$ and $n+1$, the subset $V_{+}^{(n)}\subset M_{+}^{\prime}$ can be isotoped down to the negative boundary $M^{\prime}$ of $W_{cap}$, and the same is true for $N(V_{+}^{(n)})\simeq V_{+}\subset M_{+}^{\prime}$. The final claim follows from the fact that it is obtained by performing a certain number, say $k$, of contact surgeries on $OB(V_{+},\operatorname{Id})$, which can at most kill $k$ generators in the homology group of the corresponding degree. ∎
$V_{+}\times\mathbb{S}^{1}$
$V_{-}\times\mathbb{S}^{1}$
$V_{-}\times D^{2}$
surgery handles
$M_{+}^{\prime}$
$N\times D^{*}\mathbb{S}^{1}$
[…]
3. An algebraic perspective on tightness
By the work of Bourgeois and van-Koert [BvK10], the contact homology, defined rigorously on the homology level by Pardon [Par19] and Bao-Honda [BH23], vanishes for overtwisted manifolds. Motivated from this, Bourgeois and Niederkrüger [BN10] introduced the notion of algebraically overtwisted manifolds as follows.
Definition 3.1.
Let $Y$ be a contact manifold and $\operatorname{CH}(Y;\Lambda)$ the contact homology of $Y$ over the Novikov field $\Lambda$. One says that:
  1. (1)
    $Y$ is algebraically overtwisted if $\operatorname{CH}(Y;\Lambda)=0$ [BN10];
  2. (2)
    $Y$ is algebraically tight if $\operatorname{CH}(Y;\Lambda)\neq 0$.
Proposition 3.2.
We have the following properties.
  1. (1)
    If $Y$ is overtwisted, then $Y$ is algebraically overtwisted. Hence algebraically tight contact manifolds are tight.
  2. (2)
    $Y$ is algebraically overtwisted if and only if one connected component of $Y$ is algebraically overtwisted.
  3. (3)
    Let $W$ be a Liouville cobordism with convex boundary $\partial_{+}W$ and concave boundary $\partial_{-}W$. If $\partial_{+}W$ is algebraically overtwisted then so is $\partial_{-}W$.
  4. (4)
    Strongly fillable contact manifolds are algebraically tight.
  5. (5)
    The following contact manifolds are algebraically tight.
    1. (a)
      Hypertight contact manifolds;
    2. (b)
      Contact manifolds with vanishing rational first Chern class and a contact form $\alpha$, such that there is no contractible Reeb orbit of SFT degree $1$;
    3. (c)
      $1$-ADC contact manifolds.
Proof.
The first claim follows from [BvK10] and the main result of [CMP19]. Although [BvK10] uses $\mathbb{Q}$-coefficients, the equivalence with using $\Lambda$-coefficients can be found in [MZ20, Proposition 3.12]. The second claim follows from the fact that $\operatorname{CH}(Y_{1}\sqcup Y_{2})=\operatorname{CH}(Y_{1})\otimes\operatorname{CH}(Y_{2})$. The third claim follows from the functoriality of contact homology, i.e. we have a unital algebra map $\operatorname{CH}(\partial_{+}W)\to\operatorname{CH}(\partial_{-}W)$. The fourth claim follows from the fact that $\operatorname{CH}(Y;\Lambda)=0$ is an obstruction to the existence of augmentations, hence existence of strong fillings, see e.g. [MZ20, Proposition 2.11, 3.13]. Conditions in (a) and (b) of the fifth claim mean that $1\neq 0\in\operatorname{CH}(Y)$, which implies algebraic tightness by homotopy class or degree reasons. Roughly speaking, (c) is a special case of (b). (Recall that $k-$ADC contact structures have torsion $c_{1}$ by definition.) However, to cope with the asymptotic property in the definition of $1$-ADC, we can argue as follows. If $\operatorname{CH}(Y)=0$, then for any fixed contact form $\alpha$, there exists a positive number $A$, such that $1=0\in\operatorname{CH}^{<A}(Y,\alpha)$, i.e. the homology of the sub-complex generated by $\alpha$-orbits with contact action at most $A$. By the $1$-ADC condition, we can find a contact form $\alpha_{0}<\alpha$, such that all $\alpha_{0}$-Reeb orbits with action smaller than $A$ have SFT degree at least $2$. Then the functoriality $\operatorname{CH}^{<A}(Y,\alpha)\to\operatorname{CH}^{<A}(Y,\alpha_{0})$ implies that $1=0\in\operatorname{CH}^{<A}(Y,\alpha_{0})$, which is impossible due to degree reasons. ∎
Example 3.3 ($1$-ADC Bourgeois manifolds).
Starting with an open book whose binding is $1$-ADC and simply connected, according to Lemma 2.8 the associated Bourgeois contact structure is also $1$-ADC (in fact $3$-ADC), hence algebraically tight. As special cases, Bourgeois manifolds in Example 2.10 are algebraically tight. Then after contact surgeries, the contact manifold remains algebraically tight, hence tight.
Remark 3.4 (Algebraic vs. geometric).
One of the most fundamental questions in symplectic and contact topology is whether the boundary between rigidity and flexibility is captured by pseudo-holomorphic curves. In this framework, it was shown by Avdek [Avd20] that algebraic overtwistedness in dimension $3$ does not imply overtwistedness. Avdek used another tightness criterion from Heegaard-Floer homology, which is still based on pseudo-holomorphic curves. Unlike the situation in dimension $3$, in higher dimensions, it seems that algebraic tightness is the only currently available criterion to ensure tightness.
In dimension $5$, we can reinterpret the proof of [BGM22, Theorem A] to get the following.
Theorem 3.5.
$5$-dimensional Bourgeois contact structures $BO(\Sigma,\psi)$ are algebraically tight if $\Sigma$ is non-sporadic (i.e. not a sphere with $3$ or less punctures).
Proof.
The hypothesis of page non-sporadic gives that there is a strong symplectic cobordisms whose negative end consists of hypertight contact manifolds [BGM22, Theorem 3.1 and Corollary 4.1]. Now the symplectic form on this cobordism is exact (hence no sphere bubbling can occur), however the given global primitive $1$-form is not compatible with the contact structure on the negative end (this cobordism is called pseudo-Liouville in [BGM22]). In terms of functoriality in SFT, a priori this leads to deformations of the algebraic structure on the negative boundary by counting holomorphic caps in the cobordism. However, symplectic caps for this specific cobordism can be ruled out as shown in the proof of [BGM22, Lemma 5.1]. Hence the usual functoriality for contact homology holds (over Novikov coefficients), and the positive boundary is also algebraically tight, as hypertight contact manifolds are algebraically tight. ∎
By point $(3)$ of Proposition 3.2, we then have the following useful consequence:
Corollary 3.6.
Any contact manifold obtained via contact surgery on a $5$-dimensional Bourgeois contact structure $BO(\Sigma,\psi)$ for non-sporadic $\Sigma$ is algebraically tight, and hence in particular tight.
4. Obstructions to strong fillings
4.1. Homological restrictions on symplectic fillings
The main tool that we will use to obstruct strong symplectic fillings is given by the following:
Theorem 4.1.
Let $W_{0}$ be a $(2n+2)$-dimensional subcritical Weinstein domain with $c_{1}(W_{0})=0$, and $M$ any contact manifold. If $W$ is a strong filling of $M\#\partial W_{0}$, then the kernel of $H_{n}(\partial W_{0};\mathbb{Q})\to H_{n}(W_{0};\mathbb{Q})$ contains the kernel of $H_{n}(\partial W_{0};\mathbb{Q})\stackrel{{\scriptstyle 0\oplus\operatorname{Id}}}{{\hookrightarrow}}H_{n}(M\#\partial W_{0};\mathbb{Q})\to H_{n}(W;\mathbb{Q})$.
Moreover, the same conclusion holds under the hypothesis that $W_{0}$ is a flexible Weinstein domain of dimension $6$, provided that the contact structure on $M$ has vanishing $c_{1}$.
Remark 4.2.
Theorem 4.1 is stated in this way only for the purpose of constructing exotic contact structures on spheres in this paper, but in fact a stronger version of this result, giving homological information on the fillings in some other degrees for $W_{0}$ flexible, can also be obtained by similar arguments. We also point out that, when $M=\emptyset$ and $W_{0}$ is subcritical, one can use a strategy of filling by pseudo-holomorphic curves similar to [BGZ19] to obtain a similar result, which was used by the first three authors [BGM22, Theorem B] to obstruct strong fillings.
Theorem 4.1 should be understood in the context of the classification of fillings of special families of contact manifolds in higher dimensions. Taking the celebrated Eliashberg–Floer–McDuff theorem [McD91] as a starting point, many contact manifolds are expected to have unique Liouville/symplectically aspherical fillings, at least up to diffeomorphism. Up until now, results in this direction mainly apply to the contact boundary of a split manifold of the form $V\times\mathbb{D}^{2}$, up to subcritical/flexible surgeries; we refer to [OV12, BGZ19, Zho21b, Zho22a, BGM22, GKZ23].
If one considers the more general class of strong fillings, these are no longer unique, as one can take a blow-up of any given filling without affecting the boundary. However, it is plausible that all the ambiguity only comes from birational surgeries like blow-ups, which will increase the complexity of the topology of the filling; see [Zho21b, §8] for supporting evidence. Moreover, Eliashberg [Eli90] showed in dimension $3$, that if $W$ is a symplectic filling of $Y\#Y^{\prime}$, then $W$ is obtained from attaching a $1$-handle to a symplectic filling $W^{\prime}$ of $Y_{1}\sqcup Y_{2}$ (which might be connected, cf. [Bow12]). Although the higher dimensional analogue has not been established, partial generalizations were obtained by Ghiggini-Niederkrüger-Wendl [GNW16]. Incarnations of these expectations at the level of homology lead to our obstructions to strong fillings.
4.2. Symplectic Cohomology
We will approach Theorem 4.1 through the language of symplectic cohomology. More precisely, for its proof we will need both the existence of certain moduli spaces and the non-existence of others. We then start by recalling the basics of symplectic cohomology (we refer readers to [CO18] for a more detailed account) and describing in detail the moduli spaces involved.
4.2.1. Basics of symplectic cohomology
Let $W$ be a Liouville filling with strict contact boundary $(Y,\alpha)$.
  1. (1)
    For any $D>0$, which is not the period of some Reeb orbit of $\alpha$, we have a $\mathbb{Z}/2-$graded (with grading given by $\frac{\dim W}{2}-\mu_{CZ}$) filtered positive symplectic cohomology $SH_{+}^{*,<D}(W)$. Its underlying cochain complex is generated by the (hat and check versions of the) Reeb orbits with period up to $D$. We also have the filtered symplectic cohomology $SH^{*,<D}(W)$, whose cochain complex is generated by the above orbits and a Morse cochain complex of $W$. Moreover, these cohomologies fit into a tautological long exact sequence
    $\ldots\to H^{*}(W)\to SH^{*,<D}(W)\to SH_{+}^{*,<D}(W)\stackrel{{\scriptstyle\delta}}{{\to}}H^{*}(W)[1]\to\ldots$
  2. (2)
    We define $\delta_{\partial}:SH_{+}^{*,<D}(W)\to H^{*}(Y)[1]$ to be the composition of the connecting map $\delta$ and the restriction map in cohomology associated to the inclusion $Y\hookrightarrow W$. Then, according to [Zho21b, §3.1], $\delta_{\partial}$ can be defined directly by counting rigid configurations consisting of a holomorphic plane in $W$ and a tail of gradient flow line in $Y$, with respect to any auxiliary Morse function on $Y$. When $\delta_{\partial}(x)$ has a non-trivial component in $H^{d}(Y)$, then we have
    $d-\frac{1}{2}\dim W+\mu_{CZ}(x)-1=0,$
    where $\mu_{CZ}(x)$ is the Conley-Zehnder index computed with respect to the trivialization induced by the holomorphic plane which contributes to said non-trivial component in $H^{d}(Y)$.
  3. (3)
    One can also define symplectic cohomology for strong fillings [^2] The existence of positive symplectic cohomology for strong fillings is a consequence of the asymptotic behaviour lemma [CO18, Lemma 2.3]. with coefficient the Novikov field over $\mathbb{Q}$. Regarding the analytical foundation, one needs to apply virtual techniques (e.g. polyfolds, virtual fundamental cycles, Kuranishi techniques) to define symplectic cohomology for general strong fillings going beyond semi-positive symplectic fillings. For our purposes, we can use the construction of Hamiltonian-Floer cohomology for general symplectic manifolds in [Par16], and the results in this paper rely only on the fact that a virtual count/perturbed count can be arranged to be the geometric count when transversality holds, e.g. by [Par19, Proposition 4.33].
4.2.2. Neck-stretching
We will repeatedly apply neck-stretching [${}^{+}$03] to moduli spaces in symplectic cohomology. Roughly speaking, this has two uses:
  1. (1)
    exclude certain moduli spaces, by action/index considerations;
  2. (2)
    establish the existence of certain other moduli spaces by comparison with the situation in the “standard filling”.
Both uses will yield a simpler presentation of differentials as well as of other structural maps (e.g. $\delta_{\partial}$) in symplectic cohomology. Here we present a very brief account on how to apply neck-stretching to moduli spaces in symplectic cohomology, and refer readers to [CO18] (see also [Zho21b, Zho22a]) for details.
Let $(Y,\alpha)$ be a contact type hypersurface in a symplectic filling $W$. Assume that $Y$ divides $W$ into the union of a cobordism $X$ from $Y$ to $\partial W$ and a filling $W^{\prime}$ of $Y$. For any almost complex structure $J$, which is compatible with the contact structure near $Y$, we can find a $[0,1)-$family of compatible almost complex structures on $\widehat{W}$ starting at $J_{0}=J$ and converging, for $t\to 1$, to the almost complex structures on the completions $\widehat{X}$, $\widehat{W^{\prime}}$ and $Y\times\mathbb{R}_{+}$ obtained by stretching $J$ in the Liouville vector direction along $Y$. Holomorphic curves and/or Floer cylinders for this family of almost complex structures $J_{t}$, will then converge to SFT buildings for $t\to 1$, in such a way that the top-level curve, i.e. the curve contained in $\widehat{X}$ will develop negative punctures asymptotic to Reeb orbits on $Y$. Let $\Gamma$ be the set of Reeb orbits to which these negative punctures are asymptotic. There are then mainly two constraints on the top-level curve:
  1. (1)
    Action constraint:
    $\mathcal{A}_{H}(\text{positive end})-\mathcal{A}_{H}(\text{negative end})-\sum_{\gamma\in\Gamma}\int\gamma^{*}\alpha>0$
    where $\mathcal{A}_{H}$ is the symplectic action [Zho21b, (2.1)] for the Hamiltonian $H$, which is well-defined if $\widehat{X}$ is exact.
  2. (2)
    Index constraint: the top-level curve must have a non-negative expected dimension. More precisely, when $c_{1}(X)$ is torsion, then
    $m-\sum_{\gamma\in\Gamma}\left(\mu_{CZ}(\gamma)+\frac{\dim W}{2}-3\right)\geq 0,$
    where $m$ is the expected dimension of the moduli space without negative punctures and $\mu_{CZ}(\gamma)+\frac{\dim W}{2}-3$ is the SFT degree of $\gamma$. Note that when we speak of expected dimension we will always consider the unparametrised moduli space where we quotient out any translation symmetries.
We now want to explicitly describe the moduli spaces that we are reduced to count when applying neck-stretching to the contact boundary of a symplectic manifold (recall the discussion in (3) of Section 4.2), seen inside the completion of the latter.
For this purpose, let $Y$ be a contact manifold equipped with a non-degenerate contact form $\alpha$. Let also $H$ be a (time-dependent) Hamiltonian on $\widehat{Y}:=\mathbb{R}_{t}\times Y\simeq(\mathbb{R}_{+})_{r}\times Y$, where $r=e^{t}$, such that $H$ is 00 on $(0,1]_{r}\times Y$ and is the same as the standard Hamiltonian with slope $D$ on the cylindrical end as in the usual definition of symplectic cohomology. The almost complex structure $J$ is also assumed to be cylindrical in the cylindrical end, as in the usual symplectic cohomology setup. We will refer to such a pair $(H,J)$ as admissible.
In this setting, given a multiset $\Gamma=\{\gamma_{1},\ldots,\gamma_{k}\}$ of good Reeb orbits and $x,y$ non-constant Hamiltonian orbits, we can consider the compactification ¯$\overline{{\mathcal{M}}}_{Y,H}(x,y,\Gamma)$ of the moduli space
${\mathcal{M}}_{Y,H}(x,y,\Gamma)={\raisebox{5.0pt}{$\left\{u:\mathbb{R}_{s}\times S^{1}_{t}\backslash\{p_{1},\ldots,p_{k}\}\to\widehat{Y}\;\left|\begin{array}[]{c}\partial_{s}u+J(\partial_{t}u-X_{H})=0,\;\lim_{s\to\infty}u(s,\cdot)=x,\\ \lim_{s\to-\infty}u(s,\cdot)=y,\;\lim_{p_{i}}u=\gamma_{i}\end{array}\right.\right\}$}\left/\raisebox{-5.0pt}{$\mathbb{R}$}\right.}\,.$
Here, the notation $\lim_{p_{i}}u=\gamma_{i}$ means that $u$ is asymptotic to $\gamma_{i}$ at $p_{i}$ viewed as a negative puncture with a free asymptotic marker; recall that in this region the Hamiltonian is zero and the equation is the usual Cauchy-Riemann equation. Note that the compactification ¯$\overline{{\mathcal{M}}}_{Y,H}(x,y,\Gamma)$ is a mixture of Floer-type breaking at non-constant Hamiltonian orbits, as no breaking can happen at a constant orbit of $H$ by symplectic action reasons, and of SFT building breaking at the lower level. This is the moduli space that may appear after applying neck-stretching to the moduli space defining the differential for $SH^{*}_{+}$.
Similarly, for a given singular chain $C\subset Y$ and a fixed $0<\eta\ll 1$, we consider the compactification ¯$\overline{{\mathcal{M}}}_{Y,H}(x,C,\Gamma)$ of the following moduli space:
${\mathcal{M}}_{Y,H}(x,C,\Gamma)={\raisebox{5.0pt}{$\left\{u:\mathbb{C}\backslash\{p_{1},\ldots,p_{k}\}\to\widehat{Y}=\mathbb{R}_{+}\times Y\left|\begin{array}[]{c}\partial_{s}u+J(\partial_{t}u-X_{H})=0,\;\,p_{i}\neq 0,\\ \displaystyle\lim_{s\to\infty}u(s,\cdot)=x,\;\,\lim_{p_{i}}u=\gamma_{i},\\ u(0)\in\{1-\eta\}\times C,\end{array}\right.\right\}$}\left/\raisebox{-5.0pt}{$\mathbb{R}$}\right.}\,,$
In actual constructions, to avoid working with the infinite-dimensional model $C^{*}(Y)$ of singular chains, one can use a finite-dimensional one: for instance, one can just use simplicial complexes from a triangulation or Morse complexes thanks to the choice of an auxiliary Morse function on $Y$, in which case $C$ is replaced by the stable manifolds of the gradient flow. This is the moduli space that may appear after applying neck-stretching to the moduli space defining $\delta_{\partial}$.
Proposition 4.3.
Let $(Y,\xi)$ be the contact manifold $M\#\partial W_{0}$ in Theorem 4.1 with the additional assumption that $M$ is Liouville fillable, and $0\neq\beta\in\operatorname{image}(H^{n}(W_{0};\mathbb{Q})\to H^{n}(\partial W_{0};\mathbb{Q}))$. Then there exist a contact form $\alpha$ for $\xi$, an admissible pair $(H,J)$ of a Hamiltonian and an almost complex structure, a Morse function $f$ on $Y$, a $\mathbb{Q}$-linear combination $\sum a_{i}x_{i}$ of non-constant Hamiltonian orbits of $H$ (of Conley-Zehnder index $2$ when it makes sense, i.e. $c_{1}(M\#\partial W_{0})=0$), and a $\mathbb{Q}$-linear combination $\sum b_{i}p_{i}$ of index $n$ critical points of $f$, such that the following properties are satisfied:
  1. (1)
    $\sum b_{i}p_{i}$ represents $0\oplus\beta$ in $H^{n}(M;\mathbb{Q})\oplus H^{n}(\partial W_{0};\mathbb{Q})=H^{n}(Y;\mathbb{Q})$ via Morse cohomology of $f$;
  2. (2)
    $\sum a_{i}\,\#\overline{{\mathcal{M}}}_{Y,H}(x_{i},y)=0$, where $y$ is a non-constant Hamiltonian orbit; to be precise, by this we mean that the moduli spaces of the type ¯$\overline{{\mathcal{M}}}_{Y,H}(x_{i},y)$ that are of expected dimension at most 00 are cut out transversely, and the sum $\sum a_{i}\,\#\overline{{\mathcal{M}}}_{Y,H}(x_{i},y)$ for those with expected dimension 00 vanishes;
  3. (3)
    ¯$\overline{{\mathcal{M}}}_{Y,H}(x_{i},y,\Gamma)=\emptyset$ for all $i$, $y$ and $\Gamma\neq\emptyset$;
  4. (4)
    ¯$\overline{{\mathcal{M}}}_{Y,H}(x_{i},S_{p_{i}},\Gamma)=\emptyset$ for $\Gamma\neq\emptyset$ and $\sum_{i,j}a_{i}b_{j}\,\#\overline{{\mathcal{M}}}_{Y,H}(x_{i},S_{p_{j}},\emptyset)\neq 0$, where $S_{p}$ denotes the stable manifold of $p$ w.r.t. $\nabla f$.
Proof.
For the reader’s convenience and because the two situation require different arguments, we divide the proof in two cases. First, we look at the setting where $W_{0}$ is flexible (of dimension $22n+$ 2) and $c_{1}$ of the contact structure at the contact connected sum vanishes. Here, we use index arguments to understand whether the moduli spaces considered are (generically) empty. Then, we deal with the case of $W_{0}$ subcritical and no assumption on $c_{1}$ of the contact structure on $M$. In this case, the Conley-Zehnder index is (potentially) not well defined, so we argue instead using topological and energy arguments thanks to the specific topological splitting of the form $W_{0}=V\times\mathbb{D}^{2}$ given by [Cie02b].
Case 1: $W_{0}$ is flexible Weinstein (of dimension $22n+$ 2) and $c_{1}(M)=0$.
By [BEE12, MS18], we have that $SH^{*}_{+}(W_{0})\to H^{*}(W_{0})[1]$ is an isomorphism for the standard flexible filling $W_{0}$. Moreover, according to [Laz20a], there exists a contact form $\alpha$ on $\partial W_{0}$ such that all Reeb orbits of period $<D$ are non-degenerate and have Conley-Zehnder indices at least $1$, with the additional property that $SH_{+}^{*,<D}(W_{0})\to H^{*}(W_{0})[1]$ is surjective. In the connected sum, we can assume the contact form on $M$ is large enough so that all Reeb orbits on $M$ have a period $\gg D$. Following [Laz20a, Theorem 3.15] and [Yau04], we can then assume that all the Reeb orbits of period $<D$ on the connected sum $M\#\partial W_{0}$ are those in the connected-summand $\partial W_{0}$, plus the multiple covers $\gamma_{h}^{k}$ of a Reeb orbit $\gamma_{h}$ winding around the core of the $1$-handle, for $k$ smaller than a certain $k_{0}$ depending on $D$ and the “thickness” of the contact handle realising the connected sum. Moreover, the Conley-Zehnder index of $\gamma_{h}^{k}$ is $n+2k-1$.
Let then $F$ be the Liouville filling of $M$. Using the above contact form, we have $H^{n}(W_{0})\to SH^{n,<D}(W_{0})=SH^{n,<D}(F\natural W_{0})$ is zero, where the last equality is given by [Cie02a, Fau20]. As a consequence, we have $SH^{n-1,<D}_{+}(F\natural W_{0})\to H^{n}(F\natural W_{0})$ is surjective onto $0\oplus H^{n}(W_{0})$. Therefore we have $\delta_{\partial}:SH^{n-1,<D}_{+}(F\natural W_{0})\to H^{n}(F\natural W_{0})\to H^{n}(M\#\partial W_{0})$ is surjective onto $0\oplus\operatorname{image}(H^{n}(W_{0})\to H^{n}(\partial W_{0}))$. Since $\beta\in\operatorname{image}(H^{n}(W_{0})\to H^{n}(\partial W_{0}))$, there are $x_{i}$ Hamiltonian orbits with Conley-Zehnder index $2$ and a linear combination $\sum a_{i}x_{i}$ representing an element in the positive symplectic cohomology (i.e. in particular closed w.r.t. the differential) so that $\delta_{\partial}(\sum a_{i}x_{i})=\beta$. We moreover choose a Morse function $f$ on $Y$ such that $0\oplus\beta\in H^{n}(M)\oplus H^{n}(\partial W_{0})=H^{n}(Y)$ is represented by $\sum b_{i}p_{i}$, for some index $n$ critical points $p_{i}$ of $f$.
Now we apply neck-stretching to the moduli spaces defining the differential and $\delta_{\partial}$ in order to deduce the desired properties. The expected dimension for ¯$\overline{{\mathcal{M}}}_{Y,H}(x_{i},y,\Gamma)$ is
$\mu_{CZ}(x_{i})-\mu_{CZ}(y_{i})-1-\sum_{\gamma\in\Gamma}(\mu_{CZ}(\gamma)+n-2)\leq 2-1-1-\sum_{\gamma\in\Gamma}(n-1)=-(n-1)\,\#\Gamma$
which is negative if $\Gamma\neq\emptyset$; hence (3) follows. Similarly, the expected dimension of ¯$\overline{{\mathcal{M}}}_{Y,H}(x_{i},S_{p_{i}},\Gamma)$ is
$0-\sum_{\gamma\in\Gamma}(\mu_{CZ}(\gamma)+n-2)\leq-(n-1)\,\#\Gamma\,,$
which is again negative if $\Gamma\neq\emptyset$, hence the first claim of (4) holds. When $y$ has Conley-Zehnder index larger than $1$, (2) follows from dimension counting. When $y$ has Conley-Zehnder index $1$, (2) holds as $\sum a_{i}x_{i}$ is closed for any sufficiently stretched almost complex structure by [Zho21b, Lemma 2.15] and there are no other degenerations in the fully stretched case by (3). For the second claim of (4), we apply neck-stretching to the moduli space contributing to $\delta_{\partial}(\sum a_{i}x_{i})$. Since we can rule out ¯$\overline{{\mathcal{M}}}_{Y,H}(y,S_{p_{i}},\Gamma)$ for any $y$ with Conley-Zehnder index at most $1$ by dimension counting, there cannot be other degeneration in the neck-stretching, and so $\sum_{i,j}a_{i}b_{j}\,\#\overline{{\mathcal{M}}}_{Y,H}(x_{i},S_{p_{j}},\emptyset)\neq 0$ corresponds to $\delta_{\partial}(\sum a_{i}x_{i})=0\oplus\beta$, which holds for any sufficiently stretched almost complex structure by [Zho21b, Proposition 2.17].
Case 2: $W_{0}$ is subcritical (and $c_{1}(M)$ is not necessarily zero).
As already pointed out at the beginning of the proof, the lack of the $c_{1}=0$ assumption for the contact structure on $M$ costs us the well-definedness of the Conley-Zehnder indices. We then argue differently, using more the topology of this situation, which allows to derive some crucial energy bounds. More precisely, according to [Cie02b], if $W_{0}$ is subcritical then it splits as a symplectic product as $W_{0}=V\times\mathbb{D}^{2}$, for some Weinstein domain $V$. This product structure and the fact that the considered almost complex structures are compatible with this splitting allows us to get the following lower energy bound for (parts of) Floer cylinders which are entirely contained in the symplectization of the piece $\partial V\times\mathbb{D}^{2}\subset\partial(V\times\mathbb{D}^{2})$, where the chosen Hamiltonian only depends on the symplectization direction $t$: the composition $\pi_{D}\circ u$, for $\pi_{D}\colon\mathbb{R}_{t}\times\partial V\times\mathbb{D}^{2}\to\mathbb{D}^{2}$ the natural projection, is a holomorphic map to $\mathbb{D}$ and we have
$\int u^{*}\mathrm{d}\lambda\geq\int(\pi_{\mathbb{D}}u)^{*}\mathrm{d}\left(\frac{r^{2}}{2\pi}\mathrm{d}\theta\right)\geq 0\,.$ (*)
The aim is then to use this energy bound (*4) to make up for the loss of the well-definedness of the Conley-Zehnder indices in this case, and still prove the desired conclusion. In order to do so, we need to carefully choose a convenient contact form for the contact structure on $M\#\partial(V\times\mathbb{D}^{2})$.
For this, we start by describing an explicit choice of contact form on $\partial(V\times\mathbb{D}^{2})$ as in [Zho22a, §2.1]. More precisely, one first rounds the corner of
$\left(\,\partial(V\times\mathbb{D}^{2})=V\times S^{1}\cup\partial V\times\mathbb{D}^{2}\,,\;\lambda_{V}+\frac{r^{2}}{2\pi}\mathrm{d}\theta\,\right)\,,$
and the desired contact form $\alpha$ on $\partial(V\times\mathbb{D}^{2})$ is then a perturbation of $\lambda_{V}+\frac{r^{2}}{2\pi}\mathrm{d}\theta$ supported on the mapping torus region $V\times S^{1}$. As explained in detail in [Zho22a, §2.1], this perturbation is achieved via an auxiliary Morse function $f$ on $V$, satisfying the following properties:
  • it is self-indexing, i.e. $f(p)>f(q)$ if and only if $\mathrm{ind}(p)>\mathrm{ind}(q)$ for every pair of critical points $p,q$;
  • $f(p)$ is approximately $1$ for every critical point $p$ of $f$;
  • using $r$ as collar coordinate given by the Liouville vector field near the boundary, $\frac{\partial}{\partial r}(\frac{1}{f})\to\infty$ and $f\to 1$ when approaching the boundary of $V$.
Leaving out the details of the perturbation, for our purposes it is enough to point out that, for an $f$ chosen as above, the perturbed contact form $\alpha$ is given by $\lambda_{V}+\frac{r^{2}}{2\pi}\mathrm{d}\theta$ on $\partial V\times\mathbb{D}^{2}$, and by $\lambda_{V}+\frac{1}{2\pi f}\mathrm{d}\theta$ on $V\times S^{1}=\partial(V\times\mathbb{D}^{2})\,\setminus\,\partial V\times\mathbb{D}^{2}$. In particular, it satisfies the following properties.
  1. (1)
    Each critical point $p$ of $f$ corresponds to a simple non-degenerate Reeb orbit $\gamma_{p}$, which is the circle over $p$ in the region $V\times S^{1}\subset\partial(V\times\mathbb{D}^{2})$. In particular, the Reeb orbits $\gamma_{p_{i}}$’s wind around the binding $\partial V\times\{0\}\subset\partial V\times\mathbb{D}^{2}$ exactly once.
  2. (2)
    The period of $\gamma_{p}$ is $1/f(p_{i})$ (which is approximately $1$ by choice of $f$), and hence $\gamma_{p}$ has longer period than $\gamma_{q}$ if and only if $f(p)<f(q)$.
  3. (3)
    The set of all Reeb orbits with period at most $2$ is just $\{\,\gamma_{p}\;|\;p\in\mathrm{Crit}(f)\,\}$.
Now, $M\#\partial(V\times\mathbb{D}^{2})$ is obtained by contact surgery of index 00 on the disjoint union $M\sqcup\partial(V\times\mathbb{D}^{2})$, or, in other words, it is the contact boundary of a Weinstein $1$-handle attachment on the disjoint union of the Liouville filling $F$ of $M$ with $V\times\mathbb{D}^{2}$. The part of surgery (or handle attachment) in the $\partial(V\times\mathbb{D}^{2})$ component can be more precisely done inside the subset $V\times S^{1}\subset\partial(V\times\mathbb{D}^{2})$, away from every critical point of $f$ (times the $S^{1}$ factor), and more precisely in the region $f^{-1}((a,b))$, where $a$ is the minimum of $f$ and $b$ is the image via $f$ of its index $1$ critical points (recall $f$ is self indexing, so all index 00 and index $1$ critical points have same image). We use here in particular the special contact form, that we denote $\lambda$, on the connected sum given by [Laz20a, Theorem 3.15] and [Yau04] as done for Case 1. More precisely, the action of the new simple closed Reeb orbit $\gamma_{h}$ on the belt sphere of the resulting connect sum depends on the thickness of the Weinstein handle realizing this contact surgery, and we fix once and for all such thickness to be very small (which results in the action of $\gamma_{h}$ to also be very small). According to [Laz20a, Theorem 3.15] and [Yau04], one can in fact arrange that the Reeb orbits on $M\#\partial(V\times\mathbb{D})$ of period $<2$ are just the $\gamma_{p}$’s and (some of) the multiple covers of the simple Reeb orbit $\gamma_{h}$ in the belt sphere of the handle.
In our setting, we moreover claim that we can attach the $1$-handle to the $V\times S^{1}$ region in such a way that the contact form $\lambda$ satisfies the following property: whenever $p_{i}$ is not the minimum point of $f$,
$\frac{1}{f(p_{i})}-1<\int\gamma_{h}^{*}\lambda\,.$ (**)
This can easily be arranged as follows. As the quantity at the right hand side only depends on the thickness of the Weinstein handle realizing the contact surgery (which has been fixed), up to homotopy, supported in an open set $U\subset V$ so that $U\times S^{1}$ avoids the contact surgery, of $f$ among Morse functions satisfying all the properties listed above, we can move all the critical points of $f$ but the minimum to be so near to the boundary of $V$, where $f=1$, so that the desired inequality is satisfied. Note that such an open set $U$ and homotopy exist by the assumption done before that the contact surgery is done in the region $f^{-1}((a,b))\times S^{1}$, where $a$ is the global minimum of $f$ and $b$ is the image of the index $1$ critical points (recall again that $f$ is self-indexing).
Let’s now start to arrange the desired almost complex structure in order to ensure that we can indeed use the energy bound (*4) described before. On the neighborhood $\partial V\times\mathbb{D}^{2}$ of the binding, the contact structure is given by
$\xi_{\partial V}\oplus\langle\,\partial_{x}+yR_{\lambda_{V}},\partial_{y}-xR_{\lambda_{V}}\,\rangle,$
where $\xi_{\partial V}\subset T\partial V$ is the contact structure $\ker\lambda_{V}$, $R_{\lambda_{V}}$ is the Reeb vector field of $(\partial V,\lambda_{V})$ and $x,y$ are coordinates on $\mathbb{D}^{2}$. Moreover, the Reeb vector field $R$ of $\alpha$ on the neighborhood $\partial V\times\mathbb{D}^{2}$ of the binding is just $R_{\lambda_{V}}$. We then use an almost complex structure $J$ on $\mathbb{R}_{t}\times\partial V\times\mathbb{D}^{2}$ satisfying the following properties:
$J:\xi_{\partial V}\to\xi_{\partial V}\text{ is compatible with }\mathrm{d}\lambda_{V},\quad J(\partial_{x}+yR_{\lambda_{V}})=\partial_{y}-xR_{\lambda_{V}}\,\text{ and }\,J(R_{\lambda_{V}})=-\partial_{t}\,.$
Note that with this choice of $J$, the natural projection $\pi_{D}$ from $\mathbb{R}_{t}\times\partial V\times\mathbb{D}^{2}$ to $\mathbb{D}^{2}$ is $(J,i)$-holomorphic, where $i$ is the standard complex structure on $\mathbb{D}^{2}$. In particular, we can indeed use the energy lower bound described in (*4).
Let’s now describe the Hamiltonian used. First on the filling $W_{0}=V\times\mathbb{D}^{2}$ of $\partial(V\times\mathbb{D}^{2})$ (i.e. before connect sum with $M$), one can consider a Hamiltonian $H$ with slope $2$ which vanishes on the filling. There are then two non-constant Hamiltonian orbits $\check{\gamma}_{p},\hat{\gamma}_{p}$ from each one of the previously described Reeb orbit $\gamma_{p}$’s with $p$ critical point of the Morse function $f$. Then, by [Zho22a, Proposition 2.6], for every
$\beta\in\operatorname{image}(H^{n}(V\times\mathbb{D})\to H^{n}(\partial(V\times\mathbb{D})))\simeq H^{n}(V)\,,$
there exists $\sum a_{i}\check{\gamma}_{p_{i}}$ in $SH^{*}_{+}$ and $\delta_{\partial}(\sum a_{i}\check{\gamma}_{p_{i}})=\beta$ with $\mathrm{ind}(p_{i})=n$; in fact, one simply has $\beta=\sum a_{i}p_{i}$.
We now look at the connected sum with $M$ described before, and use the special contact form $\lambda$ on $M\#\partial(V\times\mathbb{D}^{2})$ which has been previously described. Recall in particular that the Reeb orbits on $M\#\partial(V\times\mathbb{D})$ of period $<2$ are just the $\gamma_{p}$’s and (some of) the multiple covers of the simple Reeb orbit $\gamma_{h}$ in the belt sphere of the handle, and that the inequality (**4) holds for any $p_{i}$ which is not the minimum point of $f$.
Note that, since the $p_{i}$’s in the above identity $\beta=\sum a_{i}p_{i}$ all have index $n$, the associated $\gamma_{p_{i}}$’s have the minimum period among all the Reeb orbits of the form $\gamma_{p}$. In particular, (2), (3) and the first claim of (4) will directly follow from action considerations, provided we can rule out the possibility of the $\gamma_{h}^{k}$’s, or the corresponding Hamiltonian orbits, appearing as negative asymptotics. To do this, we use an argument inspired from [CDvK16, Lemma 5.5], which leverages the energy lower bound described in (*4) in the binding region. (In fact, our situation is simpler than that in [CDvK16], as we are working with the trivial open book $\partial(V\times\mathbb{D})$.)
As the only Reeb orbits of action $<D$ on the connect sum have been arranged to be the ones from $\partial(V\times D^{2})$ and the iterates of the Reeb orbit on the belt sphere of the handle, for the moduli spaces in the conclusion of the proposition, it is enough to look at $\overline{\gamma}_{h}^{k}\in\{\check{\gamma}_{h}^{k},\hat{\gamma}_{h}^{k}\}$ and $\Gamma$ be a multiset of Reeb orbits taken from the $\gamma_{h}^{k}$’s. Then, for any curve $u$ in ¯$\overline{{\mathcal{M}}}_{Y,H}(\check{\gamma}_{p_{i}},\overline{\gamma}_{h}^{k},\Gamma)$, we have
$\int_{S^{1}}\check{\gamma}_{p_{i}}^{*}\lambda-\int_{S^{1}}(\gamma_{h}^{k})^{*}\lambda-\sum_{\gamma\in\Gamma}\int_{S^{1}}\gamma^{*}\lambda\;\geq\;\int_{u^{-1}(\mathbb{R}_{t}\times\partial V\times\mathbb{D}^{2})}u^{*}\mathrm{d}\lambda\;\geq\;\int_{u^{-1}(\mathbb{R}_{t}\times\partial V\times\mathbb{D}^{2})}(\pi_{\mathbb{D}}u)^{*}\mathrm{d}(\frac{r^{2}}{2\pi}\mathrm{d}\theta)=1\,,$
where the second inequality follows from (*4), and the last equality follows from the fact the fact that the linking number of the $\gamma_{p_{i}}$’s and of the $\gamma_{h}^{k}$’s around the binding $\partial V$ are $1$ and 00 respectively. Hence we arrive at a contradiction with the inequality (**4), provided such a curve $u$ in ¯$\overline{{\mathcal{M}}}_{Y,H}(\check{\gamma}_{p_{i}},\overline{\gamma}_{h}^{k},\Gamma)$ exists. This establishes (2), (3) and the first claim of (4).
Lastly, using the Viterbo transfer map we see that, in the connected sum, $\sum a_{i}\check{\gamma}_{p_{i}}$ on $M\#\partial(V\times\mathbb{D})$ is mapped to $\sum a_{i}\check{\gamma}_{p_{i}}$ on $\partial(V\times\mathbb{D})$. Hence $\delta_{\partial}(\sum a_{i}\check{\gamma}_{p_{i}})=0\oplus\beta$ for the Liouville filling $F\natural(V\times\mathbb{D})$. Combining with the just proved emptiness of the moduli spaces of the points (2), (3) and (4) in the statement, we then get the second claim of (4) by a neck-stretching argument, as done in Case 1. ∎
Proof of Theorem 4.1.
As in the statement, assume $W$ is a strong filling of $Y$. By the universal coefficient theorem, it is equivalent to prove the dual statement that $H^{n}(W;\mathbb{Q})\to H^{n}(Y;\mathbb{Q})$ is surjective onto $0\oplus\operatorname{image}(H^{n}(W_{0};\mathbb{Q})\to H^{n}(\partial W_{0};\mathbb{Q}))$. By [CE20, Laz20b], there is a Weinstein cobordism from $M$ to $M^{\prime}$, such that $M^{\prime}$ is Weinstein fillable. Combining with [BCS15, Proposition 3.4], we can assume the Weinstein cobordism has vanishing first Chern class if $M$ does and $\dim M=5$. Then we have a Weinstein cobordism from $Y$ to $Y^{\prime}:=M^{\prime}\#\partial W_{0}$, yielding a strong filling $W^{\prime}$ of $Y^{\prime}$. It is clear that the claim follows from the surjective property for $W^{\prime},Y^{\prime}$. In other words, it suffices to prove for Liouville fillable $M$. For this purpose, let $\beta\in\operatorname{image}(H^{n}(W_{0};\mathbb{Q})\to H^{n}(\partial W_{0};\mathbb{Q}))$; we need to prove that $0\oplus\beta$ is in the image of $H^{n}(W;\mathbb{Q})\to H^{n}(Y;\mathbb{Q})$.
We will apply Proposition 4.3; let in particular $H$ and $\sum a_{i}x_{i}$ be the resulting Hamiltonian and $\mathbb{Q}$-linear combination of non-constant periodic orbits. Viewing $H$ as a Hamiltonian on $\widehat{W}$, then by applying neck-stretching along the contact boundary, we see that, for a sufficiently stretched almost complex structure, $\sum a_{i}x_{i}$ represents a class in $SH_{+}^{*}(W)$ (with Novikov field coefficient over $\mathbb{Q}$) by (2) and (3) of Proposition 4.3. Moreover, (4) of Proposition 4.3 implies that $\sum a_{i}x_{i}$ is mapped to $0\oplus\beta$ through $\delta_{\partial}\colon SH^{*}_{+}(W)\to H^{*+1}(W)\to H^{*+1}(Y)$. In particular, $0\oplus\beta$ lies in the image of $H^{*+1}(W)\to H^{*+1}(Y)$, as desired. ∎
5. Non-fillable contact structures
5.1. Tight non-fillable structures on spheres
We first construct tight, non-fillable contact structures on $\mathbb{S}^{2n+1}$, for $n\geq 2$ to prove Theorem B.
Theorem 5.1.
For every $n\geq 2$, the sphere $\mathbb{S}^{2n+1}$ admits an algebraically tight, non-strongly fillable contact structure that is homotopically standard.
Proof.
We consider the Bourgeois contact structure coming from the following choice of open book on the sphere $\mathbb{S}^{2n-1}$. For $n\geq 3$, the open book coming from the $A_{1}-$singularity, with binding given by $B=\Sigma_{n-1}(2,2,\cdots,2)\cong S^{*}\mathbb{S}^{n-1}$; this is just the positive stabilisation $\mathbb{S}^{2n-1}=OB(D^{*}\mathbb{S}^{n-1},\tau)$ of the standard open book $OB(\mathbb{D}^{2n-2},\operatorname{Id})$, where $\tau$ is the Dehn-Seidel twist. For $n=2$, an open book $OB(\Sigma,\psi)$ for the standard contact structure on $\mathbb{S}^{3}$, whose page $\Sigma$ is non-sporadic and has homology $H_{1}(\Sigma)$ of rank at least $2$; for example, one can take repeated positive stabilisations of the trivial open book.
Notice that, topologically, the Bourgeois manifold associated to these open books is of the form $\mathbb{S}^{2n-1}\times\mathbb{T}^{2}$. We then perform subcritical/flexible surgeries to topologically obtain a sphere $\mathbb{S}^{2n+1}$. More precisely, first observe that, by the $h$-principle for isotropic embeddings [EM02, Section 12.4] and for loose Legendrians [Mur12], one can always realise smooth surgeries of index $1$ and $2$ via contact surgeries. In view of this, we first perform two index $1$ surgeries that kill the generators of the fundamental group in the torus factor, to obtain a simply connected manifold with the homology of $\mathbb{S}^{2n-1}\times\mathbb{S}^{2}$. Notice that, since these generators are non-trivial in homology, the resulting cobordism will have trivial $c_{1}$. An index $2$ surgery then kills the generator in degree $2$ (which is indeed spherical by the Hurewicz theorem). In order to ensure that this surgery is possible one needs to make sure that the normal bundle is trivial: however, this is ensured by the assumption that $c_{1}$ vanishes, since the obstruction to the normal bundle being trivial is given by the second Stiefel-Whitney class $w_{2}$, to which $c_{1}$ reduces modulo $2$. Finally, the corresponding handle attachments applied to the boundary of the null-cobordism $\mathbb{D}^{2n}\times\mathbb{\mathbb{T}}^{2}$ yield a simply connected homology ball, which is then smoothly a ball by the $h$-cobordism theorem, giving that the manifold obtained by these surgeries is indeed a smoothly standard sphere. Realising these surgeries as Weinstein handle attachment according to the above mentioned $h$-principles, we thus obtain the desired contact structure $\xi_{ex}$ on $\mathbb{S}^{2n+1}$.
Now, we claim that $(\mathbb{S}^{2n+1},\xi_{ex})$ is algebraically tight. Indeed, the Bourgeois structures before surgeries are algebraically tight: for $n=2$ this is the content of Corollary 3.6, while for $n\geq 3$ these structures are in fact $1$-ADC by Example 2.10, hence algebraically tight in view of Proposition 3.2. Then, since algebraic tightness is preserved under contact surgery, the claim follows.
Notice also that the almost contact structure underlying $\xi_{ex}$ is the standard one on $\mathbb{S}^{2n+1}$, since it extends as an almost complex structure on the smooth filling which is just a ball, on which any two almost complex structures are homotopic.
To obstruct fillability, we want to apply Theorem 4.1 to the filling given by stacking the hypothetical filling of the contact structure on the sphere with the strong cobordism of Proposition 2.13. First, we remark that the $1$- surgeries can all performed along curves lying in a small neighbourhood of a fixed $\mathbb{T}^{2}$-fiber, since the isotropic $h$-principle is $C^{0}$-small and the $2$-surgery can be performed on a sphere contained in the trace of of this small neighbourhood under the initial surgery. In particular, all can be done on one side of the $\mathbb{S}^{1}$-equivariant decomposition for the Bourgeois contact structure. Now, we have $V_{+}=\Sigma\times D^{*}\mathbb{S}^{1}$ and $M^{\prime}_{+}$ is the resulting flexibly fillable contact manifold given by first capping then doing surgery as in Proposition 2.13. If $(\mathbb{S}^{2n+1},\xi_{ex})$ has a strong filling $W_{0}$, we get a strong filling $W$ of $M^{\prime}_{+}$ by stacking the strong cobordism of Proposition 2.13 on $W_{0}$. Note that there exists a homology class $\beta\in H_{n}(M^{\prime}_{+};\mathbb{Q})$ that is non-trivial in the flexible filling $W^{\prime}$ and is contained in $V_{+}\times\{pt\}$. Since we can always find a non-trivial homology class $\beta^{\prime}\in H_{n}(W^{\prime};\mathbb{Q})$ contained in $V_{+}\times\{pt\}$ by our assumption on $\Sigma$, we can take $\beta$ to be any preimage of $\beta$ under the surjective map $H_{n}(M^{\prime}_{+};\mathbb{Q})\to H_{n}(W^{\prime};\mathbb{Q})$. Since $V_{+}\times\{pt\}$ can be isotoped to $\mathbb{S}^{2n+1}$ in the cobordism, $\beta$ must be trivial in the glued strong filling contradicting Theorem 4.1 for $M=\emptyset$. ∎
Remark 5.2 (Freedom of choice of Open Book).
For concreteness we chose to take the open book determined by the $A_{1}$-singularity in the proof above. This said, any nontrivial Milnor open book whose binding is $-1$-ADC would suffice.
Theorem 5.3.
Let $(\mathbb{S}^{2n+1},\xi_{ex})$ be a exotic contact sphere in Theorem B, then $(M,\xi)\#(\mathbb{S}^{2n+1},\xi_{ex})$ is not strongly fillable for any contact manifold $M$ if $n\geq 3$. When $n=2$, the same holds as long as $c_{1}(M)$ is zero.
Proof.
Note that from the proof of Theorem 5.1, there exists a strong cobordism from $(\mathbb{S}^{2n+1},\xi_{ex})$ to a flexibly fillable contact manifold $M^{\prime}_{+}$ with vanishing first Chern class, such that there is a non-trivial homology class $\beta\in H_{n}(M^{\prime}_{+};\mathbb{Q})$ annihilated in the cobordism but not in the flexible filling. As the cobordism is obtained by blowing-down a Giroux domain and contact surgeries, which are operations performed away from a point, we get a strong cobordism from $(\mathbb{S}^{2n+1},\xi_{ex})\#M$ to $M^{\prime}_{+}\#M$ and the homology class corresponding to $\beta$ is again annihilated in the cobordism. We then apply Theorem 4.1 to obstruct fillability by stacking the blow-up cobordism on top of any hypothetical filling of $(\mathbb{S}^{2n+1},\xi_{ex})\#M$. ∎
Proof of Theorem A.
We can assume $(M,\xi)$ admits a strong filling, for otherwise, we have nothing to prove. In this case the connected sum $(M,\xi)\#(\mathbb{S}^{2n+1},\xi_{ex})$ is not strongly fillable by Theorem 5.3. Finally by Proposition 3.2, $(M,\xi)$, hence $(M,\xi)\#(\mathbb{S}^{2n+1},\xi_{ex})$ are both algebraically tight. ∎
5.2. Abundance of exotic structures on spheres
Let $(\mathbb{S}^{2n+1},\xi_{ex})$ denote a tight non-fillable contact structure constructed in Theorem 5.1. This was obtained (for dimension at least $7$) via subcritical surgery on the Bourgeois manifold $BO(D^{*}\mathbb{S}^{n-1},\tau)$, which, as originally observed by Lisi, has an associated spinal open book decomposition; c.f. [BGM22, Section 2] for details. Namely, we have a decomposition
$BO(D^{*}\mathbb{S}^{n-1},\tau)=S^{*}\mathbb{S}^{n-1}\times D^{*}\mathbb{T}^{2}\bigcup_{S^{*}\mathbb{S}^{n-1}\times\mathbb{T}^{3}}\mbox{Map}(D^{*}\mathbb{S}^{n-1},\tau)\times\mathbb{T}^{2},$
where $\mbox{Map}(D^{*}\mathbb{S}^{n-1},\tau)$ denotes the mapping torus of the Dehn-Seidel twist $\tau:D^{*}\mathbb{S}^{n-1}\to D^{*}\mathbb{S}^{n-1}$.
Lemma 5.4.
When $n\geq 3$, the contact manifold $(\mathbb{S}^{2n+1},\xi_{ex})$ is $(n-3)$-ADC.
Proof.
Lemma 2.8 gives the indices of contractible Reeb orbits on the Bourgeois manifold $BO(D^{*}\mathbb{S}^{n-1},\tau)$. On the other hand, the subcritial surgeries kill part of the topology of $BO(D^{*}\mathbb{S}^{n-1},\tau)$ making non-contractible orbits of $BO(D^{*}\mathbb{S}^{n-1},\tau)$ contractible in $(\mathbb{S}^{2n+1},\xi_{ex})$. We then need to compute the Conley-Zehnder indices of them as well. For simplicity we pick a global trivialisation of the almost contact structure underlying the Bourgeois contact structure, which is split as an almost contact structure according to Remark 2.2, by simply taking a trivialisation of $(\mathbb{S}^{2n-1},\xi_{st})$ and trivialising the tangent bundles of the $\mathbb{T}^{2}$-fibers in an equivariant manner. The Reeb orbits on $BO(D^{*}\mathbb{S}^{n-1},\tau)$ then have Conley-Zehnder indices as follows:
  1. (1)
    On $B\times D^{*}\mathbb{T}^{2}$, where $B=S^{*}\mathbb{S}^{n-1}$, the computation of the Conley-Zehnder indices is exactly the same as in Lemma 2.8. Namely, we have
    $\mu_{CZ}(\gamma_{q})=\mu_{CZ}(\gamma_{B})-2+\mbox{ind}_{q}(H)\geq\mu_{CZ}(\gamma_{B}).$
    Here, $\gamma_{q}=\gamma_{B}\times\{(q,r=0)\}$ is the contractible orbit on the binding corresponding to the critical point $q\in\mathbb{T}^{2}\times\{0\}$ of a Morse function $H$ on $D^{*}\mathbb{T}^{2}$. From the computations in [KvK16, p. 37] for the $A_{1}$-singularity, we obtain that $\mu_{CZ}(\gamma_{B})\geq n-2$, and thus
    $\mu_{CZ}(\gamma_{q})\geq n-2,$
    for every orbit in this region.
  2. (2)
    On the product $\mbox{Map}(D^{*}\mathbb{S}^{n-1},\tau)\times\mathbb{T}^{2}$, the orbits stay tangent to the pages of the open book. Moreover, for a fixed angle $\theta\in\mathbb{S}^{1}$, the closed Reeb orbits tangent to the $\theta-$page are naturally parametrized by the page $D^{*}\mathbb{S}^{n-1}$ itself, namely they are flat geodesics on $\mathbb{T}^{2}$ with rational slope $\theta$ (recall Formula (2)), and so they form an $\mathbb{S}^{1}$ Morse–Bott family when viewed as Reeb orbits on $\mathbb{T}^{3}$. As in Lemma 2.8, we locally perturb the contact form by a Morse function with a unique local maximum on $D^{*}\mathbb{S}^{n-1}$ and two critical points $q_{1},q_{2}$ lying in the zero section, respectively of index $22n-$ 2 and $n-1$. The minimal Conley-Zehnder index of the orbits on $\mathbb{T}^{3}$ after perturbing with a Morse function on $\mathbb{S}^{1}$ with critical points $p_{1},p_{2}$ is zero. By standard properties of the Conley-Zehnder index, if we denote by $\gamma_{p_{j}}$ the Reeb orbit on $\mathbb{T}^{3}$ corresponding to $p_{j}$, then the non-degenerate orbit $\gamma_{q_{i},p_{j}}$ corresponding to the pair $(q_{i},p_{j})$ has Conley-Zehnder index
    $\begin{split}\mu_{CZ}(\gamma_{q_{i},p_{j}})=&\;\mbox{ind}(q_{i})-\frac{1}{2}\dim(D^{*}\mathbb{S}^{n-1})+\mu_{CZ}(\gamma_{p_{j}})\\ =&\;\mbox{ind}(q_{i})-n+1+\mu_{CZ}(\gamma_{p_{j}})\\ \geq&\;n-1-n+1+0=0.\end{split}$
    Therefore the Conley-Zehnder index of every orbit is non-negative along this region, and their SFT degrees are at least $n-2$.
  3. (3)
    Along the interface region $S^{*}\mathbb{S}^{n-1}\times\mathbb{T}^{3}$, the situation is modelled as a smoothened corner of the product $D^{*}\mathbb{S}^{n-1}\times D^{*}\mathbb{T}^{2}$. Computations in this setting are carried out in detail in [Zho21b, Proposition 6.18 (3)], and the result is that the minimal Conley-Zehnder index is the sum of minimal Conley-Zehnder indices on $S^{*}\mathbb{S}^{n-1}$ (i.e. $n-2$) and $S^{*}\mathbb{T}^{2}$ (i.e. 00) [^3] Note that [Zho21b] considered the $S^{1}$-family of Reeb orbits, whose generalized Conley-Zehnder index has an extra $1/2$. But here we consider non-degenerate Reeb orbits after a small perturbation of the Morse-Bott family, hence the minimal Conley-Zehnder index adds.. Hence we have $\mu_{CZ}(\gamma)\geq n-2$ for every orbit $\gamma$ in this region.
  4. (4)
    The attaching of two $1$-handles and a $2$-handles create Reeb orbits of index at least $n-2$. When we attach the two $1$-handles to kill the fundamental group of $BO(D^{*}\mathbb{S}^{n-1},\tau)$, to maintain the correctness of Conley-Zehnder indices computed above, we need to choose the framing so that the induced trivialization of the contact structure (used in cases 2 and 3 above) along the isotropic circle extends to a trivialization of the core disk in the handle. In fact since the $\mathbb{T}^{2}$-fibers over the binding of the open book in the base are isotropic, these give natural choices of isotropic circles for surgery (of any slope). One then simply takes the framing of the normal bundle induced from the trivialisation on the symplectic normal induced from the projection to the base.
In total we see that after surgery all periodic orbits have degree at least $n+1-3=n-2$ (up to a certain action threshold). Thus the resulting contact structure is $(n-3)$-ADC. ∎
As proved in [Laz20a], the ADC condition is sufficient to obtain well-defined invariants of Liouville-fillable contact manifolds using positive symplectic cohomology (generated by contractible orbits) of any of its topologically simple [^4] That is $c_{1}=0$ and the inclusion of the boundary induces an injective map on $\pi_{1}$. Liouville fillings. Under a slightly stronger assumption, namely that there is a non-degenerate Reeb vector field all of whose contractible orbits have SFT degree at least $2$, Cieliebak-Oancea [CO18, Section 9.5] described a version of symplectic homology on the trivial symplectic cobordism $M\times[0,1]$ associated to the contact manifold $M$. This yields a contact invariant in view of a neck-stretching argument. In fact, for exactly the same reason as in the fillable case [Laz20a], it is enough to have $1$-ADC for this to be a well-defined invariant of the contact manifold itself. With a slight abuse of notation, we will then denote the positive symplectic cohomology of a $1$-ADC contact manifold $(M,\xi)$ simply by $SH_{+}^{*}(M)$ (i.e. $SH^{>0}_{*}(M)$ in [CO18, Section 9.5]). When $M$ is the contact boundary of a topologically simple Liouville filling $W$, we have $SH_{+}^{*}(M)=SH_{+}^{*}(W)$.
Proposition 5.5.
When $n\geq 5$, there exists a fillable contact structure $(\mathbb{S}^{2n+1},\xi^{\prime})$ homotopic to the standard almost contact structure, which is $1$-ADC, and such that $0<\dim SH^{2n-3}_{+}(\mathbb{S}^{2n+1},\xi^{\prime})<\infty$.
Proof.
We consider the Brieskorn sphere $Y:=\Sigma(n,\ldots,n,n+1,p)$, where $p$ is a prime number with $p\gg n$. By [KvK16, (14)] or [Zho22c, (5.12)], the minimal Conley-Zehnder index of a small perturbation of the standard contact form on $Y$ is $4-n$ (obtained with $|I|=n+2$, $|I_{T}|=n$, $N=1$, $T=n$ in [Zho22c, (5.12)]). The SFT degrees are then at least $2$. The minimal Conley-Zehnder indices of other Morse-Bott families of Reeb orbits of the natural contact form (after suitable perturbation) is $6-n$ (given by $N=2$, $T=n$ in [Zho22c, (5.12)]) as $p\gg n$, and so they have SFT degree $4$. As a consequence, we have that $Y$ is $1$-ADC. Moreover, by the Morse-Bott spectral sequence and the index gap above, we have $\dim SH_{+}^{(n+1)-(4-n)}(Y)=\dim SH_{+}^{2n-3}(Y)=1$, solely contributed to by the family with the minimal Conley-Zehnder index. Now, by [KvK16, Proposition 3.6], $Y$ is homeomorphic to a sphere. Then there exists $M$, such that the $M$-th iterated self connected sum #$\#^{M}Y$ is diffeomorphic to the standard sphere. Since $\dim Y\geq 11$, the contact sum preserves the $1$-ADC property and $\dim SH_{+}^{2n-3}(\#^{M}Y)=\dim\bigoplus^{M}SH^{2n-3}_{+}(Y)=M$ by [CO18, Theorem 7.1, Proposition 9.19]. To obtain the standard almost contact structure, we choose $Y^{\prime}$ to be the flexibly fillable sphere with the opposite homotopy class  [^5] To be precise, here “opposite” is to be interpreted as the inverse w.r.t. the natural group structure on the space of almost contact structures on the sphere given by the connected sum operation. of almost contact structures as that of #$\#^{M}Y$, as in [Laz20a, Corollary 1.6]. Then $Y^{\prime}\#^{M}Y$ is a standard sphere with the standard almost contact structure, and it is $1$-ADC. Moreover, $SH^{2n-3}_{+}(Y^{\prime})=0$ when $n\geq 5$. Hence we may take $Y^{\prime}\#^{M}Y$ as the desired contact sphere. ∎
We now have all the needed ingredients to prove Theorem D.
Proof of Theorem D.
One first realises the given almost contact structure as the contact boundary $(M,\xi_{flex})$ of a flexible Weinstein domain using Eliashberg’s h-principle [CE12]. For $i\in\mathbb{N}$, we consider the iterated connected sums $Y_{i}:=(\mathbb{S}^{2n+1},\xi_{ex})\#(M,\xi_{flex})\#^{i}(\mathbb{S}^{2n+1},\xi^{\prime})$ with the contact sphere $(\mathbb{S}^{2n+1},\xi_{ex})$ from Theorem 5.1 and $(\mathbb{S}^{2n+1},\xi^{\prime})$ from Proposition 5.5. Notice that $M$ is $1$-ADC, as it is flexibly fillable [Laz20a], and that $SH^{*}_{+}(M)$ is supported in degree smaller than $n+1$, in view of the vanishing of symplectic homology for flexible Weinstein domains. Since $SH_{+}^{*}(\mathbb{S}^{2n+1},\xi_{ex})$ is supported in degrees at most $n+1$ by Lemma 5.4, then if $12n-3>n+$ 1, i.e. $n\geq 5$, we have $\dim SH^{2n-3}_{+}(Y_{i})=N\cdot i$, where $N=\dim SH^{2n-3}_{+}(\mathbb{S}^{2n+1},\xi^{\prime})$, by [CO18, Theorem 7.1, Proposition 9.19]. As a consequence, $Y_{i}$ are pairwise distinguished by $SH^{*}_{+}$. They are tight and have no strong filling by the same arguments as in the proof of Theorem A. ∎
Remark 5.6.
When $n=3$, $(\mathbb{S}^{2n+1},\xi_{ex})$ is not $1$-ADC, which costs us the well-definedness of the positive symplectic cohomology. When $n=4$, it is not easy to find a degree such that positive symplectic cohomology is different for varying summands in the connected sum. In those cases, we can use positive symplectic cohomology for augmentations in [Zho23, §3.2.3] [^6] The positive symplectic cohomology of $1$-ADC manifolds is the special case for the trivial augmentation., which was motivated by the work of Bourgeois-Oancea [BO09]. Then we can obtain a “contact invariant” by numerating through all possible augmentations. However, the current status of the contact homology [Par19, BH23] can not deduce that such contact invariant is independent of various choices. Assume the invariance of such theory, one can upgrade Theorem D to cover $n=3,4$.
6. Liouville fillable contact spheres without Weinstein fillings
We now explain how to use surgery techniques to deduce the existence of Liouville fillable contact structures that are not Weinstein fillable on spheres. This requires a refinement of the surgery arguments used above and we begin with the following:
Lemma 6.1.
Let $(M^{2n+1},\xi)$ be a contact manifold of dimension at least $5$ whose underlying almost contact structure is stably trivial (as an almost contact structure). Then there is an almost complex bordism $W$ from $(M,\xi)$ to a homotopy sphere $\Sigma$ (with some almost contact structure).
Proof.
This follows immediately from results in [BCS14], building on work of Kreck [Kre99]. For the reader’s convenience we describe the steps in the proof.
First since the contact structure is stably parallelizable it follows that the tangent bundle of $M$ is stably trivial (forgetting the complex structure) so that the classifying map of its tangent bundle factors through a point:
[…]
By surgery as in [Kre99, Lemma 2] we can add handles to $M\times I$ to inductively kill homotopy groups up to dimension $n$; let $W$ be the corresponding cobordism. Notice that the positive boundary of $W$ is a homotopy sphere by construction.
We now claim that the classifying map of the stable tangent bundle of $W$ is constant as well. This can be argued as follows. First, the stable triviality of the tangent bundle of $M$ ensures that the normal bundles of the embedded spheres on which we do surgery are again stably trivial. Because the codimensions of the attaching spheres are here $n+1\geq 3$, this is moreover equivalent to actual triviality. Lastly, one can choose framings in such a way that the stable trivialisation of the tangent bundle extends over the handles to all of $W$ and the claim then follows.
Thus, since $TW$ is stably trivial as a real bundle, we can endow a stabilization of it with an almost complex structure, i.e. we found a stable almost complex structure on $W$. The conclusion of the lemma then directly follows from the fact that, according to [BCS14, Lemma 2.16], one can destabilize the stable almost complex structure on $W$ that we just found to a genuine almost complex structure which agrees (up to homotopy) with the given one on the negative end. ∎
As we already seen in Lemma 5.4, it is important to keep track of the Conley-Zehnder indices for non-contractible orbits to understand the indices of the Reeb orbits that become contractible on the spheres obtained via the flexible surgeries. For this, assume the contact manifold $Y$ has vanishing rational first Chern class. If we fix a complex trivialisation $\tau$ of the contact structure (or sum of copies of the contact structure, if the first Chern class in non-zero) over the $2$-skeleton, or equivalently a trivialisation of the complex determinant bundle associated to the contact structure, then we can speak of the Conley-Zehnder indices for any orbits with respect to $\tau$, which we will denote by $CZ_{\tau}(\gamma)$. One can then define notions of index positivity and ADC, that take into account all periodic orbits and not just those that are contractible. In this case we will say that $(Y,\tau)$ is generalized ADC following [Zho22b, Definition 3.1].
Proposition 6.2.
Suppose that $V$ is Liouville with $c_{1}(V)$ torsion, then the contact boundary $\partial(V\times\mathbb{D}^{2})$ is generalized ADC with respect to any trivialization $\tau$.
Proof.
This proposition was essentially established in [Zho21b, Theorem 6.3]; we give some further details here for the readers’ ease.
We first choose a trivialisation $\tau^{\prime}_{0}$ of $\det_{\mathbb{C}}\oplus^{N}TV$ for some $N\in\mathbb{N}_{+}$ over $V$, which induces a trivialisation $\tau_{0}$ of $\det_{\mathbb{C}}\oplus^{N}T(V\times\mathbb{D}^{2})$ over $V\times\mathbb{D}^{2}$. Following the proof of [Zho21b, Theorem 6.3] and the notations therein, $\partial(V\times\mathbb{D}^{2})$ can then be decomposed into $Y^{2}_{\rho,f,g}$, diffeomorphic to $V\times S^{1}$, and $Y^{1}_{\rho,f,g}$, diffeomorphic to $\partial V\times\mathbb{D}^{2}$. The Reeb orbits (with period up to some threshold) on $Y^{2}_{\rho,f,g}$ are contractible and have positive SFT degrees. The Reeb orbits on $Y^{1}_{\rho,f,g}$ are described in [Zho21b, Proposition 6.7], where the Conley-Zehnder indices w.r.t. $\tau_{0}$ are computed using the Conley-Zehnder indices of the $\partial V$-component w.r.t. $\tau^{\prime}_{0}$. In particular, those orbits have arbitrarily high Conley-Zehnder indices w.r.t. $\tau_{0}$.
Lastly, for any trivialisation $\tau$ of $\det_{\mathbb{C}}\oplus^{N}T(V\times\mathbb{D}^{2})|_{\partial(V\times\mathbb{D}^{2})}$, the difference from the Conley-Zehnder index of $\gamma$ w.r.t. $\tau$ and $\tau_{0}$ is given by $\langle\,\tau-\tau_{0},\gamma\,\rangle$, where we view $\tau-\tau_{0}$ as a class in $H^{1}(\partial(V\times\mathbb{D}^{2}))$. Then the proof of [Zho21b, Theorem 6.3] implies that $\partial(V\times\mathbb{D}^{2})$ is generalized ADC w.r.t. $\tau$. ∎
We now prove Theorem C from the Introduction.
Proof of Theorem C.
We let $V$ be a Liouville domain as constructed by Massot-Niederkrüger-Wendl [MNW13], so that $V\cong N\times I$ and $N^{2n-1}=G/\Gamma$ is a quotient of a solvable Lie group by a co-compact lattice. More precisely, the contact structures on the boundary components of $V$ come from left invariant structures on $G$, and are hence parallelizable. The same holds for the (symplectic) tangent bundle on $V$, and thus for $V\times\mathbb{D}^{2}$.
Now set $M=\partial(V\times\mathbb{D}^{2})$ to be the contact boundary of the product. The underlying almost contact structure is stably trivial and we can apply Lemma 6.1 to obtain an almost Weinstein cobordism $W$ to some homotopy sphere. Combining Proposition 6.2 with [Laz20a], and realising $W$ as a flexible Weinstein cobordism using [CE12], we obtain an ADC contact structure on the convex end, which is a homotopy sphere $(\Sigma,\xi_{\Sigma})$. Moreover, this has a natural Liouville filling obtained by attaching Weinstein handles to $V\times\mathbb{D}^{2}$, which in particular has trivial symplectic cohomology by [Cie02b, BEE12]. The cohomology of the resulting filling is non-trivial in degree $12n-$ 1 corresponding to the fundamental class of $N$.
Using the ADC property and the vanishing of its symplectic cohomology, according to [Zho21b, Corollary B] the cohomology of any filling of $(\Sigma,\xi_{\Sigma})$ with vanishing first Chern class, and in particular any Weinstein filling, has cohomology identical to the cohomology of the natural Liouville filling, and hence has non-trivial cohomology in degree $12n-$ 1, which is bigger than $n+1$ under our assumption of $n\geq 3$. So we deduce that $(\Sigma,\xi_{\Sigma})$ is Liouville, but not Weinstein, fillable.
Now since the set of smooth, oriented homotopy spheres with the operation of connect sum forms a finite group, up to taking self connected sums we can assume that $\Sigma\cong\mathbb{S}^{2n+1}$ is diffeomorphic to the standard sphere, which then also has an ADC contact structure with a Liouville filling with vanishing symplectic cohomology that is homologically not Weinstein. In other words, we have found a contact structure on the standard sphere which is Liouville but not Weinstein fillable.
To get the standard almost contact structure we use the fact that all stably trivial almost contact structures on a standard sphere can be realised via contact structures (cf. [BCS14, Lemma 2.17]) which are Weinstein fillable. Taking connected sum of these with the previously constructed Liouville non-Weinstein fillable structures then gives one structure as in the statement of Theorem C. ∎
One can now easily get infinitely many distinct contact structures as in Theorem C by taking the connected sum with the infinitely many examples of homotopically standard and flexibly fillable contact structures on spheres from [Laz20a]. Arguing mutatis mutandis we also obtain Theorem E from the Introduction. We also point out that the $5$-dimensional case remains open since, for all known examples of four-dimensional Liouville domains $V$ with disconnected boundary, the manifold $V\times\mathbb{D}^{2}$ is Weinstein by Breen-Christian [BC21].